### v1

# Embodying Control in Soft Multistable Grippers from morphofunctional co-design

Abstract
Soft robots are distinguished by their ability to adapt to various environments and interaction forces through shape-changing and responsiveness to external stimuli. Their flexible materials provide inherent safety, allowing them to perform tasks that are nearly impossible for rigid robots. However, controlling the configuration of soft robots is challenging due to their highly nonlinear material response and infinite deflection degrees of freedom, often necessitating sophisticated sensors and complex closed-loop control systems. One potential solution is to discretize the infinite-dimensional configuration space of soft robots into a finite but sufficiently large number of functional shapes. This approach facilitates modeling and control without sacrificing compliance and adaptability. This study explores a co-design strategy for a class of pneumatically actuated soft grippers with multiple encoded stable states, enabling desired functional shape and stiffness reconfiguration. Combining a PneuNet bending actuator with dome-shaped bistable units allowed us to encode stable configurations and design them to exhibit desired kinematic and stiffness responses. This architecture enables position and force regulation of soft robots without closed-loop control. We present an energy-based analytical model for these soft multistable grippers based on a lattice of lumped parameter elements. This model maps the soft robots’ infinite-dimensional configuration space into discrete stable states, allowing us to predict the system’s final state and the dynamic behavior between each state upon actuation. Our approach introduces a general method to capture the soft robots’ response with the lattice lumped parameters using automatic relevance determination regression, facilitating inverse co-design. The resulting computationally efficient model enables us to explore the configuration space in a tractable manner, allowing the inverse co-design of our robots by setting desired targeted positions with optimized stiffness of the set targets. Our model is validated against FE simulations and experimental tests for different target points, demonstrating the potential of discretizing the modeling and design of soft robots using multistable structures. We showcase the utility of our modeling framework through inverse design for various applications and tasks, combining diverse mechanical behaviors into the same soft robot. Our strategy of mapping the modeling and co-design of soft grippers into a discretized space offers a framework for controlling soft robots by exploiting the nonlinear mechanics of multistable structures, thus opening avenues for embodying mechanical intelligence into soft structures.
Keywords Modeling  $\cdot$ embodied control  $\cdot$ mechanical intelligence  $\cdot$ multistability  $\cdot$ inverse design  $\cdot$ soft robotics  $\cdot$ embodied intelligence
1 Introduction
Soft robots have gained interest due to their ability to interact with their environment, adapt to external stimuli, and protect against external disturbances [1, 2, 3]. The inherent safety provided by the characteristic low modulus of materials used in soft robots allows them to perform nearly impossible tasks for their rigid counterpart[4]. The interplay between soft mechanics and controls intrinsic to soft robotics gives rise to innovative solutions for tasks ranging from simple grasping to complex manipulation and interaction[5, 6, 7].However, this interplay also poses challenges complicating their modeling and control. These challenges include their infinite dimensionality, material nonlinearity, and large deformations that most soft robots exhibit[8]. As a result, sensory systems[9] and complex models/algorithms [10, 11] are required to implement closed-loop control, leading to computationally demanding models to represent the robot’s behavior. Embedding sensing and control in the robots’ architecture [12] offers a new strategy to address some of the main problems of soft robotics reducing the use of complex algorithms and models to predict the robot’s behavior. For example, granular material-based universal grippers [13] can adapt to different object shapes and grasp them without the need for closed-loop control based on their mechanical response. Zoe et al. [14] developed a methodology to use various soft actuators as pneumatic sensors, enabling distance sensing, shape sensing, profile scanning, and stiffness determination. By incorporating the robot’s geometry and body into the design, these examples illustrate the possibility of closing the control loop to realize desired tasks by the intelligent co-design of the robot’s morphology and functional behavior.
Multistable structures offer an alternative path to achieve control of soft robotics with feedback via the programming of input-specific stable and defined shapes[15, 16, 17]. These types of structures display more than one energetically favorable configuration, giving the possibility to reach different final shapes from specific actuation inputs. They are often the result of assembling multiple classical bistable substructures including constrained beams/trusses [18, 19], constrained dielectric elastomers [20], shells [21, 22, 23, 24], compliant mechanisms [25], or inflatable structures [26]. Combining multistable structures with soft robots can reduce the number of degrees of freedom and complexity to achieve control in soft robotic devices, ultimately enhancing their functionality by providing the ability to reach desired set points without closed-loop feedback[27, 28, 29]. Presently, a wide variety of soft machines leverage mechanical instabilities to improve the design and performance of their soft machines [30, 31, 32, 33] by programming different stable configurations and utilizing the rapid energy release during snap-through instability [34, 35, 36, 37]. Conrad et al. show an electronics-free controller for a complex pneumatic system [38] based on soft logic gates that can be incorporated into different soft machines. Similarly, Peretz et al. utilized mechanical instabilities to create fluidic logic and reproduce valving functionality using entirely soft elements[39]. Melancon et al. [40] used a bistable origami pattern to program different target points to its actuator, achieved by a single pressure source and predefined pressure path. Van Raemdonck et al. [41] geometrically tuned an actuator to generate complex actuation sequences from a single input by delaying the unit snap-thought. Moreover, by leveraging the stable points in the structure, the robot’s final state can be predicted based solely on its structural response, thereby defining set points in the mechanical behavior of the structure. This approach can simplify the complexity of predictive or data-driven algorithms currently needed to design and control soft robots, ultimately making soft robots more accessible and practical. However, the full potential of using multistability for soft robotics can only be realized by establishing computationally efficient methodologies for their morphological and functional co-design that circumvent the long runtime roadblocks imposed by the predominant use of conventional Finite Element (FE) packages to predict the geometry and response characteristics associated with the desired stable states.
We introduce a framework for the modeling, control, and inverse design of soft robotics by mapping into a discretized space the desired kinematic configurations leveraging multistability (see Fig. 1). We achieve this by incorporating different dome-shape programmable units addressable via pneumatic actuation featuring bistable, pseudo-bistable, and monostable [21, 42] mechanical behaviors (see Fig. S2). As a result, different set points in the form of specific kinematical shapes are embodied into the system’s response (see Fig 1), giving our soft robots the inherent capability of reaching different positions with zero error. This characteristic allows for robust soft robot design, as the bistable units provide intrinsic kinematic control dictated by the embodied mechanical response. The position control inherent to the designed multistable structure can serve as set points for triggering different control actions. The discretized response afforded by the proposed topology enables the introduction of an energy-based analytical modeling framework, allowing the efficient shape and structural property prediction of a multistable soft robot appendage topology (Dome Phalanx Finger - DPF) that ultimately enables the inverse co-design of the robot’s functionality and control via the morphology. Our modeling approach is based on lumped parameter elements that can capture different mechanical behaviors of the dome unit via mechanics-informed identification, mapping the geometry of our robot to its mechanical response. The collective interaction of these elastic elements dictates the soft robot’s configuration, stiffness, and dynamic response. We demonstrate the versatility of our approach by performing the inverse co-design of different DPF topologies, giving specific target positions and geometrical constraints while maximizing the programmed states’ stiffnesses. Finally, we show that the time response can be incorporated into the DPF morphology by leveraging the viscous effects of the material, giving a new avenue for programming this class of robots’ time response, illustrated through pick-and-place applications. The presented modeling and inverse co-design approach leverages geometry and the resulting structural response to encode an electronics-free form of mechanical intelligence that simplifies the actuation and control of soft robots with implications on their accessibility, recyclability, and cost.
Figure 1: Multistable Soft Robot with Pre-programmed Position and Stiffness. The four-finger robot demonstrates four grasping set points and embodied control. The geometry and mechanical behavior of the robot encode the input and output behaviors in the energy landscape (G(s)). Feedback H(s) is derived from the dynamic response of the robot, which yields different attractors for each stable set point in the robot. The system’s output shows four set points as static minima in the energy landscape and attractors in the phase portrait. Set points (energy minima) are accessed by activating the dome units: State 1 $\rightarrow$ Stress-free state and State 2 $\rightarrow$ Inverted state.
2 Results
2.1 Embodied position and stiffness control
We start by incorporating dome-shaped units into the PneuNet bending actuator topology[43] (see Fig. S1) to create our Dome Phalanx Finger (DPF) topology (Fig. 2a). The finger derives its behavior from the mechanical response of the dome-shaped shell elements and the interaction between each segment. As the units invert, the top section of the finger expands, while the limiting layer (see Fig. S1d) retains its original length. This results in global curvature and different final stable states, depending on the number of active units (see Fig. S2c).
These dome-shaped units program different mechanical responses and stable states into the system, effectively discretizing the infinite-dimensional deformation space into a manageable number of kinematic configurations dependent on the unit cell geometry, each attainable via well-defined open-loop inputs. The mechanical response and stable configurations of the DPF can be tuned by adjusting the dome height per thickness ($H_{i}$), dome thickness ($t$), number of segments, unit cell size (UC), limiting layer thickness ($t_{lim}$), unit cell length ($U_{L}$), air chamber dimensions ($W_{c}$ and $L_{c}$), and the spacing between adjacent cells ($U_{sep}$) (see Fig. 2a and Fig. S1a for reference). The dome unit geometric parameters ($H$ and $t$) determine whether the mechanical response is monostable (Fig. S1c (i)), metastable (Fig. S1c (ii)), or bistable (Fig. S1c (iii)). The other geometrical parameters influence the interaction between neighboring segments without altering the unit’s stability, while, affecting the global curvature.
Given the broad range of geometrical configurations, each finger can be geometrically tuned to exhibit different mechanical responses: a bistable response, where the final shape is retained after dome inversion[21, 22, 44] (Fig. 2b $H_{i}$ = 4 mm and $H_{i}$ = 5 mm); a pseudo-bistable (metastable) response, where each unit undergoes a snap-through instability and the shape is retained for a time $t=\tau$, before returning to its stress-free state from the combined effects of the unit’s viscoelastic material behavior and dome geometry [34, 45, 46] (Fig. 2b, H = 3 mm); and finally, a monostable response, where the material behavior dominates over the geometric response and the finger responds similarly to a PneuNet bending actuator.
Figure 2: Dome Phalanx Finger (DPF) with different encoded mechanical responses and behaviors. a) 5 segment DPF geometry parameters. ($i=5$) b) Effect of dome height on finger response (Every segment with the same height). $H_{i}$ = 5 mm and $H_{i}$ = 4 mm show a bistable behavior, and $H_{i}$ = 3 mm shows a pseudo-bistable response (self-resetting system). c) Time lapse of two gripper architectures showing both bistable and pseudo-bistable responses.
These unique properties can be further exploited in a gripper architecture. By changing the dome height, one can program a gripper composed of two fingers to pick and hold an object after actuation (Fig. 2c, H = 4 mm) or pick an object and release it as the units reset (Fig. 2c, H = 3 mm), the latter of which is ideal for pick-and-place applications typically performed by conventional soft grippers. The number of stable states that can be encoded using the DPF morphology equals that of the number of constitutive units where each unit is tunable independently. Consequently, different unit geometries can be combined to achieve multiple stable states or combined behaviors within the same overall geometry. This allows for different energy minima and time responses to be embodied within the same finger, which can be accessed by a single pressure input and serve as discrete and functional robot configurations. To effectively incorporate, combine, and predict the finger’s behavior based on its geometry, a simple yet reliable modeling framework is necessary. Given the ample design space and the computational cost of using finite element (FE) models for our finger topology, we propose an energy-based spring model that captures the complex mechanical behaviors needed to design multistable soft robots with desired functional characteristics.
2.2 Model derivation and mechanics-informed parameter determination
Establishing an analytical model to capture the global behavior of the DPF requires consideration of both the unit cells’ and the global geometric parameters. By capturing the contribution of each constitutive unit of the DPF and their interactions, we can construct an energy landscape with a simplified model, where each of the minima corresponds to a stable state and a final position of the system. In this way, the state of the system ($q$) can be extracted from the resulting strain energy potential via a minimization process, which dictates the programmed stable shapes with their geometrical and stiffness characteristics.
Figure 3: An energy-based model for static and dynamic analysis of multistable soft robots based on the DPF topology. a) Spring lattice model with linear, rotational, and nonlinear springs. b) Static stable states predicted by the lattice model for $H_{1}$ = $H_{2}$ = 4 mm, $H_{3}$ = $H_{4}$ = 4.5 mm, and $H_{5}$ = 5 mm. c) Dynamic response for the soft robot with three bistable units and two metastable units ($H_{1}=4$ mm, $H_{2}=4.5$ mm, $H_{3}=5$ mm and $H_{4}=H_{5}=3$ mm). The model captures the snap-through of each bistable unit and the snap-back of the pseudo-bistable units.
The DPF geometry is represented by a lattice array comprising nonlinear, linear, and torsional springs for the static response and point masses and dash-pots for the dynamic response (Fig. 3a). The energy for the springs used in the lattice can be written as follows[47].
Nonlinear: $\text{{Nonlinear:} }E_{NL}=\frac{1}{2}k_{b}x^{2}\left(1+\left((1-\alpha)\left(\left(\frac{x}{d}\right)^{2}-2\frac{x}{d}\right)\right)\right)$ (1)
Linear: $\text{{Linear:} }E_{L}=\frac{1}{2}k_{l}\cdot x^{2}$ (2)
Torsional: $\text{{Torsional:} }E_{T}=\frac{1}{2}k_{\vartheta}\cdot\vartheta^{2}.$ (3)
The individual stiffness and spring connectivity allow us to capture the limiting layer and dome inversion effects. We use a nonlinear spring to represent different dome mechanics (Fig. S1c), which are controlled by the parameters $k_{b}$, $\alpha$, and $d$. Linear and rotational springs are utilized to capture the limiting layer of the finger by combining stretching and bending energy (see Fig. 3a). The connection between the limiting layer and the nonlinear springs is modeled as an infinitely rigid connection, given that the strain is negligible compared to the rest of the lattice elements. By minimizing the sum of all contributions to the system ($E_{tot}=\sum_{i=1}^{n}E_{L}+E_{NL}+E_{T}$), the static problem is solved, and all possible finger configurations are predicted (see Fig. 3b).
One of the primary challenges of spring-lattice models is accurately mapping the geometric parameters of the system to its stiffness constants. To address this, we integrate Finite Element (FE) simulations with Automatic Relevance Determination (ARD) regression [48] to derive expressions for $k_{b}$, $\alpha$, and $d$—the nonlinear spring parameters in Eq. 1—as functions of dome height ($H$), dome thickness ($t$), and dome curvature ($R$) (see Fig. S1). We considered different dimensionless relations ($\pi_{i}$) based on the most relevant geometric interactions affecting the dome’s energy [21]. These mechanics-informed interactions enhance the generality of the analysis by linking the mechanics to the unit’s geometry. For our model, we assume all spring constants can be expressed as functions of the shell’s load-carrying capacity ($\pi_{1}=\frac{t}{H}$), curvature-to-thickness ratio ($\pi_{2}=\frac{t}{R}$), and dome shallowness ($\frac{H}{R}$). Consequently, each parameter can be formulated as:
[…]
Where $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ is a matrix of all the possible candidates and interactions of the non-dimensional relations, as shown in Eq. S.12. By using ARD, we can automatically determine the relevance of each feature in our model by examining the values of the hyperparameters, which determine the relevance of the weights $\xi_{\alpha}$, $\xi_{k}$, and $\xi_{d}$. As a result, we derive a general equation in terms of the three non-dimensional parameters (Eq. S.14, Eq. S.15, and Eq. S.16, see section A.2.3 for details). Using these derived expressions and a gradient-based optimization algorithm (see section A.2.4 for details), the final shape of the finger can be predicted by minimizing the system’s energy, which is now a function of the geometric parameters of the structure.
A comparison between FE simulations and our model for different stable states of the finger (see Fig. 3b) shows good agreement with the energy-based lattice model. The model is further extended to capture the structure’s dynamic response, allowing us to observe the material’s viscoelastic behavior and analyze metastable cases that energy minimization alone cannot predict. The dynamic response for different unit cell geometries is shown in Fig. S5b, and for different five-segmented fingers in Fig. S5c. The energy minimization approach and the dynamic model predict the same final state of the system after reaching a steady state. To explore the combined behavior of different dome geometries, we modeled a five-segment DPF topology with bistable and metastable units (see Fig. 3c). We stacked three bistable units with different dome heights ($H_{1}=4$ mm, $H_{2}=4.5$ mm, and $H_{3}=5$ mm) and two metastable units with the same dome height ($H_{4}=H_{5}=3$ mm). This configuration allows us to achieve three unique set points in the form of stable states and a gripper state that holds its shape for a given predictable time $\tau$ (see Fig. 3c). As shown in Fig. 3c, three different snap-through events and two snap-backs (dome-unit resets) are captured, adequately representing the mechanics of the combined phenomena.
2.3 Position and stiffness inverse design
Based on the model’s performance and its agreement with the FE simulations, we pose an inverse problem to co-design different DPF topologies capable of reaching a desired pre-programmed final position (embodied position control). Since the discrete states of the system are already determined by the stable states, the optimization algorithm only searches within the fully actuated configurations, making it computationally efficient. We utilized Bayesian optimization [49] to iterate over various numbers of segments, dome heights for each of the segments ($H_{i}$), unit separation ($\text{U}_{\text{sep}}$), and unit length ($\text{U}_{\text{L}}$) (see Fig. 2a for parameter reference). Two different objective functions are utilized: one based on position and targeting a specific tip coordinate (Position Design), and the second one targeting a specific tip coordinate while maximizing the finger’s stiffness (Position + Stiffness Design) (see section A.3).
Position Design: Given a target tip coordinate [x,y] ($\text{Target}_{xy}$), the inverse problem objective function can be written as:
$\min_{H_{i},t,\text{Unit}_{L},\text{Unit}_{s}}\sqrt{\left(\text{Target}_{xy}-\text{Tip dis}\right)^{2}}$ (5)
[…]
where the Tip dis represents the tip point shown in Fig. 2a (Tip), which coincides with the top node of the finger’s last unit. A domain constraint is imposed to guarantee that every dome unit is higher than the one behind ($H_{i+1}>H_{i}$), which in certain cases can lead to up to N possible stable states (N been the number of segments).
Five different tip coordinate targets were tested to evaluate the performance of our model and the optimization algorithm. For each case, the coordinates in $x$ and $y$ were given to the algorithm (Target Position in Fig. 4a), and the optimization was run by incrementally adding one unit per run and calculating the objective function (Eq. 5). To reduce computational cost, a maximum of 8 units was defined (Fig. 4b), although more units can be added if the objective requires a longer finger geometry. The geometry with the minimum objective function value (Obj in Table S1) was selected and plotted to test if the desired position was achieved (see Fig. 4a). Geometric parameters, coordinate targets, and objective function values are reported in Table S1. The final stable state and its comparison with each target can be observed in Fig. S6. To validate our optimization algorithm, we 3D printed the five different geometries shown in Table S4 using NinjaFlex 85A and measured the final tip position after dome inversion. An error of less than 15% was measured for each sample, showing good agreement with the optimization algorithm and the lattice model.
The topologies generated by this methodology exhibit deterministic behavior, meaning that kinematic states are consistently attained as long as the gripper’s geometric parameters remain unchanged (see Fig. 4c and Movie 2). Specifically, we cycle each gripper topology between the initial and fully activated states, verifying the gripper’s ability to achieve the same aperture and positioning across multiple cycles (Tests 1-3 in Fig. 4c). This characteristic enables robust robot design, as the bistable elements provide intrinsic kinematic control dictated by the gripper’s mechanical response. Notice that the pressure for activating the units is not controlled using sensors and closed-loop control, illustrating the embodied self-regulation from the designed multistability. Consequently, the position control inherent to the designed multistable structure serves as control set points that can be embodied in the robot’s geometry and achieved with a single, open-loop controlled actuator.
Figure 4: DPF inverse design and tests a) Results for Target 1 and 3 predicted by the model after the optimization. b) Objective function vs number of segments. The minimum value is selected for every target position. c) The gripper attains different stable kinematical states that can be predicted with the proposed energy-based model and are transferred to 3D-printed samples. d) Restriction domain $R_{c}$ for stiffness optimization of DPF. e) Experimental comparison between position and position+stiffness optimization.
Position + stiffness design: The finger’s stiffness is calculated by computing the work necessary to perturb it from a given (programmed) stable configuration. We determine the system’s total energy by solving:
$\Pi=\sum_{i=1}^{n}\left[E_{L}(q)+E_{NL}(q)+E_{T}(q)\right]-W(q),$ (6)
where the external work $W$ requires prior knowledge of the direction of the external force and its path. Given this limitation, we utilized an alternative approach by creating a circular domain of radius $R_{c}$, within which none of the nodes of our lattice model are permitted to enter. By imposing this domain constraint, we establish a method to determine the system’s state at a distance of $R_{c}$ from the fully activated stable state. To implement this alternative approach, we impose an optimization constraint using Lagrange multipliers [50], given by:
$\min_{q}\left(\sum_{i=1}^{n}E_{L}(q)+E_{NL}(q)+E_{T}(q)\right)-\lambda g(x(q))$ (7)
$g(x(q))=\max(0,R_{c}-||x(q)-C||)$
where $R_{c}$ is the radius and $C$ is the center of the circle, respectively. By determining the system’s energy as a function of $R_{c}$ (see Fig. 4d), we can maximize it to increase the system’s stiffness. Using this methodology, the geometry of the DPF can be optimized to target a specific position with maximum stiffness in the neighborhood of that stable state (see Fig. 4e). We can rewrite Eq. 5 as a weighted optimization given by:
$\min_{H_{i},\text{Unit}_{L},\text{Unit}_{s}}w_{1}\sqrt{\left(\text{Target}_{xy}-\text{Tip dis}\right)^{2}}+w_{2}\frac{1}{\bar{\Pi}\left(R\right)}$ (8)
[…]
where $w_{1}$ and $w_{2}$ are the weights assigned to each objective, and $\bar{\Pi}(R)$ is the average energy for a range of $R_{c}$ values. The number of segments (N) is kept consistent with the results obtained from the previous optimization (Eq. 5), meaning that only $H_{i}$, $\text{Unit}_{L}$, and $\text{Unit}_{s}$ are considered to maximize stiffness. The results of the stiffness and position optimization are shown in Table S2. Moreover, we experimentally test Target 5 geometries produced by both optimization procedures (see Fig. S9b). By following this approach, we increased the stiffness of the Target 5 geometry by approximately 40% (see Fig. 4e), which consequently enhances the achievable grasping force.
2.4 Dome Phalanx Robot (DPR) performance
Figure 5: Dome Phalanx Robot (DPR) assembly and applications: Fingers 1 to 4 ($A_{1}$ $A_{4}$) can be changed for different applications and mechanical behaviors. a) Multistable robot with four different set points: (i) Stable state 1 with 0 inverted units, (ii) Stable state 2 with two inverted units, (iii) Stable state 3 with four inverted units, and (iv) Stable state 4 with six inverted units. b) $A_{1}$ and $A_{3}$ are optimized for position and stiffness, while $A_{2}$ and $A_{4}$ are optimized for position. $A_{2}$ and $A_{4}$ exert $F_{Dis}$, and $A_{1}$ and $A_{3}$ exert $F_{stt}$ ($F_{stt}>F_{Dis}$). c) Bistable + Metastable finger behavior: Strain energy and pressure vs. time response. d) Snapshots of embodied pick-and-place tasks: A single pressure input is given for the robot to pick the object (Commanded response) and then release it depending on the viscoelastic material response (Programmed response)(see Movie 5).
To explore the capabilities of our design approach and the advantages of incorporating instabilities in soft robotics, we fabricated a soft robot (Dome Phalanx Robot - DPR) by combining four DPFs ($A_{1}$ to $A_{4}$), as shown in Fig. 5a. We assembled the four DPFs using 3D printing a rigid base that prevents interaction between the fingers, imposes appropriate boundary conditions considered by our model, and couples our robot to a UR3 robotic arm (see Movie 1). This robot can be assembled by either combining fingers with the same topology ($A_{1}=A_{2}=A_{3}=A_{4}$), resulting in uniform global functionality, or by using two sets of actuators ($A_{1}=A_{3}$ and $A_{2}=A_{4}$), which provides multiple functionalities within the same robot body.
We demonstrate three applications using the proposed design methodology and FDM 3D printing to create multiple robot topologies. Specifically, we focus on three different applications: 1) A multistable robot with at least three set points/stable states (see Fig. 5a); 2) A configuration with two finger pairs that can exert different grasping forces but reach the same final position (see Fig. 5b); 3) A combination of bistable and metastable dome geometries (see Fig. 2c and d), where the viscoelastic response of the material allows us to encode pick and place robotic behavior into the robot geometry (see Fig. 5c and d).
1) Multistable soft robot: We further leverage the characteristics of our multistable DPF by programming different set points within the same geometry. Using the optimization algorithm, we can program three stable states by minimizing for three discrete configurations (2, 4, and 6 active units). This design approach can be implemented by modifying the objective function as follows:
$\min_{H_{i},\text{Unit}_{L},\text{Unit}_{s}}w_{j}\sqrt{\left(\text{Target j}_{xy}-\text{Tip dis j}\right)^{2}}+w_{M+1}\frac{1}{\bar{\Pi}\left(R\right)}$ (9)
[…]
where $M$ is the number of targeted stable states, and $j$ is the stable state number ($0<j\leq M$). By doing this, we can optimize for three target positions and maximum stiffness in the fully activated state (six inverted units). For this application, we impose the number of set points, the target tip position for each set point, and the number of units. While the model could be expanded to further modify the unit cell’s dimensions and determine the optimal number of units, for simplicity, we fixed the number of units at six to ensure at least two active domes per stable state. We tested our system by performing an inverse design with three different target positions (see section A.3.3 for reference), using the same methodology described previously (section 2.3). An average error of 12% was found between the desired positions and the three final stable states, as shown in Fig. 5a ((i) - (iv)).
This methodology demonstrates the versatility of our DPF and provides an alternative approach to encoding multiple set points within the finger’s body by modulating its unit cell geometry. The local geometric modifications create more accessible stable states, all of which can be controlled using a single, open-loop driven pressure inlet per finger. Since the curvature and stiffness of the system are functions of the number of active units (see Fig. S9b), the grasping force and carrying load capacity of the DPR also become functions of the active domes. Consequently, the robot can also program different set points for force (see Movie 3).
2) Multiple functional robot: For the second application, we utilized the results shown in section 2.3, where two different objective functions were considered to create two different finger topologies that achieve the same tip position but exert different grasping forces. Using the data from Table S1 and Table S2, we designed a robot arrangement where two fingers ($A_{1}$ and $A_{3}$) produce more grasping force than the others ($A_{2}$ and $A_{4}$). We tested the grasping force using a compliant mechanism composed of two bars that compress as the grasping force increases (see Fig. 5b). We characterized the mechanism’s force vs. aperture to measure the robot’s grasping force (see Movie 4). Our experiments show that the two different objective functions result in a force application of 1.18 times ($F_{Dis}$ and $F_{stt}$) while reaching the same position in their fully activated stable state.
3) Embodied soft robotic tasks:For the third application, we incorporate the time variable into our system. By leveraging the viscoelastic response and geometry of our dome structures, we can pre-program the duration for which the structure remains in its activated state before returning to its original state. As demonstrated in section 2.1, we can modulate the geometry to create a fully metastable finger (see Fig. 2c). However, the same finger can exhibit both bistable and metastable behaviors by combining bistable and metastable dome unit cells. This approach allows the finger to be encoded with a specific set point and operate in the vicinity of that point by actuating the metastable unit cells [34]. By integrating these two behaviors, we can enable task planning, such as pick-and-place operations (see Movie 5), where the robot can pick an object by activating all units, move to a specific position, and release the object without additional interventions and with just one pressure input (see Fig. 5). We tested this application by designing a five-segment finger with three bistable and two metastable units. The set point with three active units is designed with $H_{1}=4$ mm, $H_{2}=4.5$ mm, and $H_{3}=5$ mm, while the metastable units are programmed to reset after approximately 5 seconds (see section A.3.4 for details). As shown in Fig. 5c, an object can be held while all units are inverted, even when the pressure is released from the system. However, after the metastable units reset, the object falls from the gripper. Using this approach, we can program a simple pick-and-place operation as demonstrated in Fig. 5d and Movie 5.
3 Discussion
In this study, we introduced a modeling framework and inverse design methodology for soft robotics, discretizing their continuous response using multistable structures. We propose a robotic architecture, The Dome Phalanx Finger, that leverages mechanical multistability to discretize the robots’ configuration space into desired kinematics, which serve as control set points. Exploiting these characteristics, we embodied diverse static and dynamic responses, offering an alternative strategy for achieving various robot functionalities guided solely by the designed geometry and material response. Implementing an energy-based modeling approach was central to our methodology, enabling us to predict and program the DPF’s behavior accurately. Through this approach, we established a robust framework for modeling the finger’s static and dynamic responses, with a mechanics-informed parameter discovery procedure that yields high-fidelity results consistent with experimental data. Significantly, this methodology extends beyond the DPF geometry and its lattice representation, as the combination of mechanics knowledge (in this case from shell theory), data, and automatic relevance determination can be further exploited for a wide variety of geometries and reduced-order models. Utilizing the predictive capabilities of our model, we successfully implement the inverse co-design of the DPFs’ morphology and controlled functionality yielding desired target positions, accessible states, and programmable time responses based on the structure geometry and material viscoelasticity. Our results show how these characteristics can be designed to allow for the modulation of grasping force and unit reset without the need for closed-loop control. Our findings highlight the potential gains in performance when morphology and functionality are concurrently optimized (i.e., co-designed), a result only feasible via the simplified dynamics and modeling afforded by our approach. The implications of our results are far-reaching, suggesting a promising path toward developing robust soft robots. By encoding the control within the morphology and response of the system, we embody a form of mechanical intelligence offering a reduction in control complexity and forgoing the need for data-hungry models and expensive electronics. The resulting reduction in system complexity paves the way for more efficient, cost-accessible, sustainable, and adaptable soft robotics applications.
Materials and Methods
Unit cell and DPA Fabrication– The unit cells are 3D printed using fused deposition modeling (FDM) on a Raise3D Pro2 printer, utilizing Ninjatek Ninjaflex TPU filament [51]. Each finger is post-processed to correct manufacturing errors and seal the air chambers. For the DPG, we printed a PLA base that secures each of the four fingers (see Movie 1 for assembly).
Stiffness Measurement tests– The actuators were loaded onto a specially designed test rig and tested using a universal testing machine (3345 Instron) equipped with a 100N load cell (see Fig. S9b and Movie 4). The test rig constrained all degrees of freedom at the finger’s base, with six different positions equally spaced along 90° to ensure that the initial load was always perpendicular to the tested geometry. We measured the force-displacement response five times per sample at a constant velocity of 10 mm/s. Raw data was evaluated and plotted using Python 3.12.
Finite element analysis– The unit cells are modeled in Abaqus using linear elastic material properties and a combination of S3R and S4R shell elements. A dynamic implicit quasi-static approach is utilized to capture the structure’s instabilities. Geometric nonlinear analysis (Nlgeom) is used, and snap-through is triggered using a pressure load on the dome while the edges are pinned. The unit cell is initially modeled in the stress-free state. Simulations were run for different material thicknesses ($t$) and dome heights ($H$).
The full geometry of the DPA is also modeled in Abaqus using linear elastic material properties and C3D10 3D elements. Snap-through is triggered using a pressure load on every inner wall air chamber. As discussed in the analysis, a fixed boundary condition is applied to the first unit. After dome inversion, different relaxing steps are used to achieve the final stable state.
Robot actuation– Each DPF is actuated using a pressure input controlled by a ball valve. Thea, results displayed in Movies 1 - 5 are produced by applying pressure as slowly as possible to observe the multistable phenomenon.
Author Contributions
All authors designed the research. JCO and HM developed the model and derived the constants. JCO conducted FE simulations. JSR and JCO conducted the experimental measurements and model validation. All authors contributed to the writing and revision of the manuscript and supplementary materials at different stages, and AFA provided guidance throughout the research.
4 Acknowledgements
The authors gratefully acknowledge the support from the DARPA Nature as Computer (NAC) program under the guidance of Dr. J. Zhou and NSF-CAREER award No. 1944597.
Appendix A Supplementary Material
A.1 Geometry
A.1.1 Unit Cell
The system’s bistable constitutive units are composed of a dome-shaped structure encapsulated by a square chamber that allows for pneumatic actuation and reset by applying positive and negative pressure, respectably (see Fig. S1a). Each dome unit can be geometrically tuned to exhibit a bistable behavior [21], psudo-bistable(metastable) behavior where the is a snap-through instability, but the unit returns to its zero energy state[45, 46], and a monostable behavior (see Fig. S1c). These mechanical responses can be adjusted by modifying the height $H$ and thickness $t$ of the unit as shown in Fig. S1b.
Figure S1: Dome Phalanx Finger geometry and bistable unit. a) Unit cell geometry and parameters (Dome unit + air pressure chamber). b) Bistability map dependent on dome thickness ($t$) and dome height ($H$) for actuator unit cell. c) Monostable (i), Metastable (ii), and bistable (iii) dome unit behavior. d) Dome Phalanx finger geometry has two possible, stable states (Initial and fully actuated states).
A.1.2 Dome Phalanx Finger (DPF)
The DPF is multistable soft system with sequential bistable units (Fig. S1d and Fig. S2b). The finger derives its multistable behavior from bistable domed-shaped shell elements (Fig. S1a), and it can be geometrically tuned to reach and retain different final shapes after dome inversion. The DPF performance and stable kinematic configurations can be tuned by changing the dome height ($H$), Unit Cell size (UC), strain limiting thickness ($t_{L}$), air chamber dimensions ($W_{ch}$ and $U_{L}$), and the spacing between adjacent cells ($\text{U}_{\text{sep}}$). Once fully inverted, contact between adjacent units and the strain-limiting layer induces the system’s global curvature. The bistable domes on each finger segment support programmable deflections as the domes’ final positions dictate the global kinematic configuration (see Fig. S2b). The contact between the dome tip and the adjacent unit’s chamber provides an additional interaction, contributing to the final curvature after all units are activated and the pressure is removed. The strain energy stored in the gripper topology in its activated states is the summation of the individual contributions from the inverted domes (see Fig. S2c). The magnitude of the interaction between neighboring domes is also affected by dome height ($H$) and chamber thickness ($t_{c}$). Together, dome height and chamber thickness determine the degree of interaction between neighboring segments after dome inversion and yield the system’s global curvature (Fig. S2b) and the final tip position of the system (Fig. S2c).
Figure S2: Dome Phalanx Finger geometry and behavior. a) Geometrical parameters that determine the final position and strain energy of the DPF. b) Different stable states of the DPF are encoded by tunning the dome height in the unit cell. c) Strain energy (i) and Tip displacement (ii) as a function of the number of active units. Three different sets of points are encoded into the DPF.
A.2 Energy-based model and constant derivation
A.2.1 Spring Lattice array
Predicting the behavior of multistable DPFs requires consideration of both the unit cells’ and the global geometric parameters. By characterizing the contribution of each DPF subcomponent and their interactions, an energy landscape can be built where each minimum corresponds to a stable state of the system. The resulting strain from the energy minimization process dictates the programmed stable shapes with their geometrical and stiffness characteristics. We represent the DPR as a lattice comprising non-linear[47], linear, and torsional springs (Fig. S3a). The springs’ stiffness and connectivity allow us to map local extensions and rotations to defined energy contributions. We use non-linear springs featuring a ground (unstressed) state and an inverted (stressed) to capture the influence of the bistable domes on the system. To appropriately represent the bistable dome behavior, we position the non-linear springs so that the path of extension coincides with the dome’s tip position between the ground and inverted states. Linear springs capture connections between the bistable units modeled as struts. In the DPF’s case, the strain these segments experience is negligible, behaving as a strain-limiting layer. Consequently, we model these connections as nearly rigid. While the axial strains experienced by this strain-limiting layer are negligible, the layer is sufficiently thick to display bending resistance. We capture the influence of this bending stiffness by introducing torsional springs at the nodes coincident with the strain-limiting layer. The angular displacement experienced by the torsional springs is dictated by the change in the angle $\vartheta$ formed by the segments (see Fig. S3a). Minimizing the system’s total energy, $E_{tot}=\sum_{i=1}^{n}E_{L}+E_{NL}+E_{T}$ where $n$ is the number of units, constrains the space of possible interactions for the lattice elements into a discrete set of stable states. The rich configuration space of our DPF requires establishing a rational method for providing the initial guess and initializing the minimization process. To this end, we implement a geometric base model to generate initial guesses for the optimization algorithm so that the obtained states are in the neighborhood of physically feasible configurations with improved computational time. For this model, we define $k_{l}=\frac{\sqrt{3}}{2}Et_{lim}$ and $k_{t}=\frac{2}{\sqrt{3}}\frac{Et_{lim}}{12(1-\nu^{2})}$ Where $E$ is the young modulus, $\nu$ is the Poisson ratio, and $t_{lim}$ is the thickness of the limiting layer shown in Fig. S2a.
A.2.2 Material Visco-elastic response
Material time response is captured using a variation of the Kelvin-Voigt viscoelastic constitutive model[52]. The model consists of two components in parallel: a spring (representing the elastic behavior) and a dashpot (representing the viscous behavior) see Fig. 3a.
Non-linear Spring (Elastic Element): The spring represents the material’s dome behavior, as explained before. The element follows Eq. 1.
Dashpot (Viscous Element): The dashpot represents the material’s viscous behavior, meaning it resists deformation by dissipating energy as heat. The force exerted by the dashpot is proportional to the velocity of deformation. Mathematically, this relationship is described by $F=\eta\frac{dx}{dt}$, where $\eta$ coefficient is obtained by fitting the dynamic behavior to the experiments for different loading conditions. We obtained $\eta=0.2$ s-1 and utilized it for our dynamic analysis.
A.2.3 Parameter Tuning
To determine each of our lumped model’s parameters, the DPF’s fundamental unit cell is isolated into two different parts (see Fig. S3a). The bistable unit fitted as a nonlinear spring following Eq. 1 and the strain limiting layer represented by a liner spring to capture stretching and two rotational springs to capture the bending rigidity. Each of the constants for these springs needs to be tuned for different geometrical cases to produce an accurate model for inverse co-design. Given Eq. 1, the parameters to be determined for the nonlinear spring are $k_{b}$, $\alpha$, and $d$. These parameters are functions of the dome unit’s mechanical behavior (stiffness and energy barrier) as well as the interaction between the dome and the wall of the air chamber of the subsequent unit cell.
The procedure to find each constant as a function of the geometrical parameters of the dome unit cell is as follows. First, different finite element simulations of the unit cell are utilized to determine the best-fitted values for $k_{b}$, $\alpha$, and $d$ for different geometrical cases ($H$ and $t$, same as show in Fig. S1). Eq. 1 is fitted to the FE data using the nonlinear least square method to obtain the overall behavior of these parameters in terms of $H$ and $t$. Results for these fitted parameters can be observed in Fig. S3b, and the energy-displacement curve with the fitted parameters can be observed in Fig. S4a.
Figure S3: a) DPF fundamental unit cell isolated into different lumped parameter spring elements. (Non-linear + linear + rotational springs) b) Non-linear spring parameter behavior over different dome heights $H$ and dome thickness $t$.
Once the values are obtained, we use Automatic Relevance Determination(ARD) [48] regression to determine an expression for $k_{b}$, $\alpha$, and $d$ as functions of the dome height ($H$), dome thickness ($t$), and dome curvature ($R$). By using ARD, we can automatically determine the relevance of each feature in our model. Given the main task of inverse design of multistable actuators, a general and scalable model is needed to cover different design cases. We create the model by considering three different non-dimensional relations:
$\pi_{1}=\frac{t}{H}\hskip 14.22636pt\pi_{2}=\frac{t}{R}\hskip 14.22636pt\pi_{3}=\frac{H}{R}$ (S.10)
Where $\pi_{1}$ represents the load-carrying capacity of the shell, $\pi_{2}$ is the curvature-to-thickness ratio, and $\pi_{3}$ quantifies the dome shallowness. Using these three non-dimensional quantities, the following model is fitted:
[…]
$\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ is a matrix of all the possible candidates and interactions of the non-dimensional relations. The $N\times M$ matrix can be written as:
$\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})=\begin{bmatrix}\pi_{i}&\pi_{i}^{2}&\ldots&\pi^{n}_{i}&\pi_{i}\pi_{j}&\ldots&\pi^{n}_{i}\pi^{m}_{j}\end{bmatrix}$ (S.12)
where superscripts $n$ and $m$ represent the maximum degree of the polynomial use for the regression and subscripts $i$ and $j$ represent each of the non-dimensional numbers. We can fully represent this regression as:
$\begin{bmatrix}\alpha(H_{1},t_{1},R_{1})\\ \alpha_{2}(H_{2},t_{2},R_{2})\\ \vdots\\ \alpha_{k}(H_{k},t_{k},R_{k})\end{bmatrix}=\begin{bmatrix}\pi_{i}(H_{1},t_{1},R_{1})&\pi_{i}^{2}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}(H_{1},t_{1},R_{1})&\pi_{i}\pi_{j}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{1},t_{1},R_{1})\\ \pi_{i}(H_{2},t_{2},R_{2})&\pi_{i}^{2}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}(H_{2},t_{2},R_{2})&\pi_{i}\pi_{j}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{2},t_{2},R_{2})\\ \vdots\\ \pi_{i}(H_{k},t_{k},R_{k})&\pi_{i}^{2}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}(H_{k},t_{k},R_{k})&\pi_{i}\pi_{j}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{k},t_{k},R_{k})\end{bmatrix}\begin{bmatrix}\xi_{1}\\ \xi_{2}\\ \vdots\\ \xi_{k}\end{bmatrix}$ (S.13)
Note that we assume that the library of terms used to create $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ has a sufficiently rich column space that the behavior will be represented by Eq. S.11, and it can be written as a linear combination of the weights $\xi$.
The data is split into 70% train and 30% test data sets to calculate the accuracy of our model. Observation vs prediction plot and estimation error for the three different variables can be observed in Fig. S4b (i) and (ii) respectably, where a good fit between the proposed model and the test data is observed. Moreover, we determine the precision of the distributions of the weights that allow us to determine the most relevant parameters of our regression. By establishing a threshold, the model can be further reduced by removing terms until the prediction error of the original model increases by 50% its original value. This gives a reduced model with similar prediction capabilities and the most relevant geometric relations of the problem. Our reduced model is tested against all the data (train + test), where a good agreement is found even after removing the extra terms (see Fig. S4b (iii)).
Figure S4: Parameter fitting results for $\alpha$, $k_{b}$, $d$. a) Monostable (i), metastable (ii) and bistable (iii) energy-displacement model fit vs FE simulations. b) ARD regression full model observation vs. prediction plots (test data) (i), prediction error distribution for test data, and reduced model observation vs. prediction plots (all data)
The final expressions for each of the nonlinear spring parameters can be written as:
$\begin{split}k_{b}=f\left(E,\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{E}{{HR^{3}}}(-0.71975H^{4}+13.7556H^{3}t+H^{2}\left(0.511574R^{2}-9.43363Rt-18.4013t^{2}\right)\\ &+Ht\left(-1.55186R^{2}+39.3601Rt-46.6357t^{2}\right)+Rt^{2}(1.36789R-18.2071t))\end{split}$ (S.14)
$\begin{split}\alpha=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{0.0944903t^{3}}{H^{3}}+\frac{0.190736H^{2}}{R^{2}}+\frac{t^{2}(-0.256867R-0.290149t)}{H^{2}R}+\frac{H(-0.138165R-1.54765t)}{R^{2}}+\\ &\frac{0.679161t}{H}+\frac{t\left(0.705331R^{2}+0.707233Rt+3.87065t^{2}\right)}{R^{3}}\end{split}$ (S.15)
$d=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=\frac{-1.3H^{2}Rt+2.51HR^{3}-11.34HRt^{2}+464t^{4}}{R^{3}}-H-2t-\text{U}_{\text{sep}}$ (S.16)
where $H$, $t$ and $R$ are dome unit parameters, $E$ is the elastic modulus of the material and $\text{U}_{\text{sep}}$ is the separation between units (see Fig. S2).
A.2.4 Static and Dynamic Response
Using the parameters from the fitting process, we can predict stable state configurations for the DPF with the resulting model. Given that the non-linear elements of the gripper are independently stable, the expected number of states for a given dome phalanx finger topology follows $2^{n}$, where $n$ is the number of bistable segments on the finger. The model can determine the behavior of a given design in orders of magnitude less time than the FE numerical simulations, which makes it feasible for iterating through potential configurations and optimizing for the best design according to a given task (e.g., position and grasping force).
Static response: The static response of the finger is predicted by minimizing the total energy of the geometry. We performed this process using Matlab’s fmincon sqp algorithm [53]. Our algorithm can predict different final shapes depending on the initial guess geometry. Results for the model are compared with FE simulations, where various points of the limiting layer are compared (see Fig. S5a). Results show that our model can predict the final shape of the DPF for different geometric parameters with an error below 10% compared with FE simulations.
Dynamic response: The model can be further expanded by including point masses on each of the segments to represent the total mass of the system and dampers to capture the visco-elastic response of the material. By incorporating these elements and using symbolic differentiation, a dynamic model can be constructed to capture the time response of the DPF. The equations of motion for the system can be written as
$[M]\ddot{q}+[D]\dot{q}+\nabla\left(E_{L}+E_{NL}+E_{T}\right)=F_{ext}(t)$ (S.17)
Where $[M]$ is the mass matrix of the system, $[D]$ is the damping due to viscous effects, and $F_{ext}$ is the external actuation pressure. By substituting $\dot{x}_{1}=x_{2}$ and $\dot{x}_{2}=\ddot{q}$, we can rewrite the system as
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -[M]^{-1}\left(\nabla\left(E_{L}+E_{NL}+E_{T}\right)-[D]q+F_{ext}(t)\right)\end{bmatrix}$ (S.18)
The system is solved using the Runge-Kutta time integration method for loading and unloading external force conditions. Time response for different DPF geometries can be observed in Fig. S5c.
Figure S5: Static and dynamic response of the DPF. a) Stable states are found by minimizing the strain energy. A comparison with FE simulations for five domes of the same height $\rightarrow$ H = 4 mm and H = 5 mm. b) Dynamic response of lattice model with visco-elastic dampers c) Five segment DPF dynamic response for different dome H.
A.3 Geometry Optimization
A.3.1 Inverse problem $\rightarrow$ Position Optimization
The final results for the position optimization can be observed in Fig. S6, where the five target positions and the final shape predicted by the model are shown. The final geometry and objective function of the optimization algorithm are summarized in Table. S1, where we can observe that for targets 1,2 and 5, the desired position is attained with almost perfect precision.
Figure S6: Final results for inverse design problem for tip position as an objective. a) Target 1, b) Target 2, c) Target 3, d) Target 4, e) Target 5, f) Objective function vs Number of segments for the first three target positions.
Table S1: Inverse design problem results for five different targeted positions.
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.5 4.1 4.8 4.5 4.35
H2 4.7 4.3 4.8 4.7 4.35
H3 4.8 4.9 4.9 5.0 4.72
H4 5.0 4.9 5.0 4.87
H5 5.0 4.95
H6
Unit S 2.0 1.0 1.0 1.0 1.01
Unit L 9.9 7.0 7.3 7.0 8.99
UC 16.0 16.0 16.0 16.0 16.00
Obj 0.2 0.2 2.9 2.7 0.4
A.3.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
We further expand our model to include stiffness optimization in our objective function. By incorporating finger stiffness, grasping force is increased by changing the geometric parameters. Results for the optimization algorithm, where the same number of segments as in Table S1 is maintained, are shown in Table S2.
Table S2: Inverse design problem results for five different targeted positions with maximum stiffness.
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.63 4.10 4.74 4.81 4.41
H2 4.67 4.27 4.83 4.82 4.58
H3 4.90 4.49 4.85 4.87 4.71
H4 4.98 4.86 4.92 4.81
H5 4.97 4.92
H6
Unit S 2.04 1.00 1.08 1.02 1.07
Unit L 9.94 7.00 7.07 7.03 8.85
A.3.3 Inverse problem $\rightarrow$ Position + Stiffness Optimization for multiple states
We design the multistable actuator by targeting three different tip positions. These targets are intended to be reached by activating 2,4 and 6 units. The following objective function is minimized:
$\min_{H_{1},H_{2},H_{3},H_{4},H_{5},H_{6},\text{Unit}_{L},\text{Unit}_{s}}w_{1}\sqrt{\left(\text{T1}_{xy}-\text{Tip dis 1}\right)^{2}}+w_{2}\sqrt{\left(\text{T2}_{xy}-\text{Tip dis 2}\right)^{2}}+w_{3}\sqrt{\left(\text{T3}_{xy}-\text{Tip dis 3}\right)^{2}}+w_{4}\frac{1}{\bar{\Pi}\left(R\right)}\$ (S.19)
Where $w_{1}=w_{2}=w_{3}=w_{4}=1/4$ and $T_{1}$ to $T_{3}$ represent the target positions for the gripper as shown in Table S3. Results for each of the targets can be seen in Fig. S7.
Figure S7: Inverse optimization results for multistable dome phalanx actuator. a) Lattice model results for $T_{1}$(i), $T_{2}$(ii) and $T_{3}$(iii). b) Experimental results for $T_{1}$(i), $T_{2}$(ii) and $T_{3}$(iii).
Table S3: Experimental measurements for the multistable actuator.
Target Coordinates Model Experiments
State 1 ($T_{1}$) [54 , 40] [52.95 , 32.82] [ 50.2, 30.6]
State 2 ($T_{2}$) [40 , 45] [40.84 , 45.9] [ 38.2, 44.2]
State 3 ($T_{3}$) [30 , 45] [33.81 , 47.35] [ 35.2, 46.1]
A.3.4 Inverse problem $\rightarrow$ Bistable + Mestastable units
Our task planning application uses our dynamic model to calculate the resetting time of the metastable units. We fix our actuator geometry by selecting five segments with three bistable and two metastable units. The selected geometry parameters are $H_{1}=4$ mm, $H_{2}=4.5$ mm and $H_{3}=5$ mm, $\text{U}_{\text{L}}=9$ mm and $\text{U}_{\text{sep}}=1$ mm, and the last two unit heights ($H_{4}$ and $H_{5}$) is calculated to reset after 5 s with the pressure profile shown in Fig. S8a. Both dome heights are constrained to be the same ($H_{4}=H_{5}$), and the initial bistable set point can be observed in Fig. S8b. The full dynamic response predicted by the model is shown in Fig. S8c, where we can observe the three different snap-through instability and two snap-backs after 5s.
Figure S8: Pick and place task planning set points and model inputs. a) Pressure profile for task planning. b) Setpoint (stable state) for the first active three bistable units. c) Dynamic response for the bistable + metastable actuator.
A.4 Experimental Validation
To validate our inverse design approach all five proposed geometries are 3D printed using Ninjaflex 85A. FDM printing is used given its versatility for printing complex geometries from flexible materials.
A.4.1 Inverse problem experimental validation
Five geometries are tested by activating all bistable unit cells and measuring the final tip position. The final position of the finger is measured by using image processing. Results can be observed in Fig. S9a (i-v), and the target coordinates, model prediction, and experimental measurements are reported in Table S4
Table S4: Experimental measurements for five different DPFs optimized for a specific target position.
Target Coordinates Model Experiments Error
[46 , 32] [46.85 , 31.22] [42.9 , 37.7] 13%
[25 , 20] [25.34 , 20.34] [23.53 , 22.93] 10%
[25 , 40] [30.10 , 40.70] [29.4 , 43.5] 6%
[25 , 30] [28.76 , 29.87] [23.0 , 34.26] 17%
[35 , 45] [37.98 , 44.36] [36.8 , 50.77] 11%
A.4.2 Experimental validation stiffness optimization
The stiffness of each DPF is measured using an Instron universal testing machine and applying a perpendicular force to the tip of the actuator (see Fig. S9b and Movie 4). Linear behavior is observed in the neighborhood of each stable state (0, 2, 4, and 5 active units in Fig. S9b), with a different stiffness for every state. Finally, the inverse design problem results are tested by comparing the stiffness of two geometries with the same target position but two distinct objective functions (Eq. 5 and Eq. 8).
Figure S9: Experimental results for different DPF geometries. a) Printed samples for inverse design target positions (i) - (v) b) Stiffness measurement for different active units. c) Comparison between DPF geometries for position inverse design and position+stiffness inverse design for target 5.