### v2

# Embodying Control in Soft Multistable Grippers from Morphofunctional Co-design

Abstract
Soft robots are distinguished by their flexibility and adaptability, allowing them to perform nearly impossible tasks for rigid robots. However, controlling their configuration is challenging due to their nonlinear material response and infinite deflection degrees of freedom. A potential solution to these challenges is to discretize the infinite-dimensional configuration space of soft robots into a finite but sufficiently large number of functional shapes. Leveraging this approach, we present a strategy for co-designing the desired tasks and morphology of pneumatically actuated soft robots with multiple encoded stable states. Our approach introduces a general method to capture the soft robots’ response using an energy-based analytical model, the parameters of which are obtained using Recursive Feature Elimination with lasso regression. The resulting lumped-parameter model facilitates inverse co-design of the robot’s morphology and planned tasks by embodying specific dynamics upon actuation and object interactions. The resulting computationally efficient model enables us to explore the configuration space in a tractable manner, allowing the inverse co-design of our robot morphology by setting desired set points with optimized stiffness. This strategy offers a framework for simplifying the control of soft robots by exploiting the nonlinear mechanics of multistable structures and embodying mechanical intelligence into soft material systems.
Keywords Modeling  $\cdot$ embodied control  $\cdot$ mechanical intelligence  $\cdot$ multistability  $\cdot$ inverse design  $\cdot$ soft robotics  $\cdot$ embodied intelligence
1 Introduction
Soft robots have gained interest due to their ability to interact with their environment, adapt to external stimuli, and protect against external disturbances [1, 2, 3]. The inherent safety provided by the characteristic low modulus of materials used in soft robots allows them to perform nearly impossible tasks for their rigid counterpart [4]. The interplay between soft mechanics and controls intrinsic to soft robotics gives rise to innovative solutions for tasks ranging from simple grasping to complex manipulation and object interaction [5, 6, 7]. However, this interplay also poses challenges complicating their modeling and control. These challenges include their infinite dimensionality, material nonlinearity, and large deformations that most soft robots exhibit [8]. As a result, sensory systems [9] and complex models/algorithms [10, 11] are required to implement open-loop control, leading to computationally demanding models to represent the robot’s behavior. Embedding sensing and control in the robots’ architecture [12] offers a new strategy to address some of the soft robotics’ main challenges, utilizing the robot morphology to reduce the use of complex algorithms and models and relying on its mechanical behavior. For example, granular material-based universal gripper [13] can adapt to and grasp different object shapes without needing closed-loop control based on their mechanical response. Zoe et al. [14] developed a methodology to use various soft actuators as pneumatic sensors, enabling distance sensing, shape sensing, profile scanning, and stiffness determination. Moreover, mechanical instabilities have been utilized for sequential programming and fast actuation. Yang et al. [15] explore the sequential programming of a soft gripper to perform grasping and twisting motion driven solely by mechanical instabilities in the gripper’s body. Yet another example by Lue Y. et al. [16] demonstrates the use of snap-through instabilities for fast actuation and reconfiguration. These examples illustrate how the intelligent co-design of the robot’s body morphology and functional behavior enables an approach to close the control loop for realizing desired tasks.
Multistable structures offer an alternative path to achieve control of soft robotics with feedback via the programming of input-specific stable and defined shapes [17, 18, 19]. These types of structures display more than one energetically favorable configuration, giving the possibility to reach different final shapes from specific actuation inputs. They are often the result of assembling multiple classical bistable sub-structures, including constrained beams/trusses [20, 21], constrained dielectric elastomers [22], shells [23, 17, 24, 25], compliant mechanisms [26], or inflatable structures [27]. Combining multistable structures with soft robots can reduce the number of degrees of freedom and complexity to achieve control in soft systems, ultimately enhancing their functionality by providing the ability to program desired set points and sequences without closed-loop feedback [28, 29, 30], relying only on the mechanical behavior of the system [31]. Presently, a wide variety of soft machines leverage mechanical instabilities to improve the design and performance of their soft machines [32, 33, 34, 35] by programming different stable configurations and utilizing the rapid energy release during a snap-through instability [36, 37, 38, 39, 16]. Conrad et al. [40] show an electronics-free controller for a complex pneumatic system based on soft logic gates that can be incorporated into different soft machines. Similarly, Peretz et al. [41] utilized mechanical instabilities to create fluidic logic and reproduce valving functionality using entirely soft elements. Melancon et al. [42] used a bistable origami pattern to program different target points to its actuator, achieved by a single pressure source and predefined pressure path. Van Raemdonck et al. [43] geometrically tuned an actuator to generate complex actuation sequences from a single input by delaying the unit’s snap-through. Moreover, by leveraging the encoded stable states, the robot’s final configuration can be predicted based solely on its structural response. Consequently, the designed mechanical response defines the desired operational set points of the robot, all of which are stable. By transforming the infinite-dimensional deformation space into a finite number of stable states, the complexity of predicting the soft robot behavior is reduced, and its feedback can be reduced to a set of available states. This approach can simplify the complexity of predictive or data-driven algorithms currently needed to design and control soft robots, ultimately making soft robots more accessible and practical. However, the full potential of using multistability and programmed mechanical responses for soft robotics can only be realized by establishing computationally efficient methodologies for their morphological and functional co-design that circumvent the long runtime roadblocks imposed by the predominant use of conventional Finite Element (FE) packages to predict the geometry and response characteristics associated with the desired stable states.
We introduce a framework for the modeling and inverse design of soft robotics with encoded mechanical behaviors by mapping into a discretized space the desired kinematic configurations leveraging mechanical instabilities (see Figure 1). We achieve this by incorporating different dome-shape programmable units addressable via pneumatic actuation featuring bistable, pseudo-bistable (metastable), and monostable [23, 44] mechanical behaviors (see Figure S2). As a result, different set points in the form of specific kinematical shapes and dynamic responses are embodied into the system morphology (see Figure 1), giving our soft robots the inherent capability of reaching different pre-programmed positions under open-loop inputs with almost zero error. This characteristic allows for robust soft robot design, as the bistable units provide intrinsic kinematic control dictated by the embodied mechanical response. The position control inherent to the designed multistable structure can serve as set points for triggering different control actions, decoupling the actuation from the robot’s final configuration (shape). The discretized response afforded by the proposed topology enables the introduction of an energy-based analytical modeling framework, allowing the efficient shape and structural property prediction of a multistable soft robot appendage topology (in our case, the Dome Phalanx Finger - DPF) that ultimately enables the inverse co-design of the robot’s functionality and control via the morphology. Our modeling approach is based on lumped parameter elements that can capture different mechanical behaviors of the dome unit via mechanics-informed identification, mapping the geometry of our robot to its mechanical response. The collective interaction of these elastic elements dictates the soft robot’s configuration, stiffness, and dynamic response. We demonstrate the versatility of our approach by performing the inverse co-design of different DPF topologies, giving specific target positions while maximizing the programmed states’ stiffnesses. Finally, we show that different dynamic behaviors can be encoded into the DPF morphology by leveraging the viscous effects of its constitutive material. We illustrate this new avenue for programming this class of robots’ time response through pre-programmed pick-and-place tasks. The presented modeling and inverse co-design approach leverages geometry and the resulting structural response to encode an electronics-free form of mechanical intelligence that simplifies the actuation and control of soft robots with implications on their accessibility, recyclability, security, and cost.
Figure 1: Multistable soft robot with pre-programmed dynamics for positioning and stiffness. Our robot displays four grasping set points ($s_{0},s_{1},s_{3},s_{4}$) and embodied/morphological control. The robot’s morphology and mechanical behavior encodes the input and output dynamics in its energy landscape (G(s)). Feedback H(s) is derived from the dynamic response of the robot, whereby designed attractors for each stable set point are encoded in the robot geometry. The system’s output shows four set points as static minima in the energy landscape and attractors in the phase portrait, where $q$ and $\dot{q}$ are a schematic representation of the generalized coordinates of the robot. Set points (energy minima) are accessed by activating/inverting the dome units: State 1 $\rightarrow$ Stress-free state (State 1) and $\rightarrow$ Inverted state (State 2).
2 Results
2.1 Embodied position and stiffness control
We begin by modifying the common PneuNet bending actuator topology [45] via the inclusion of dome-shaped units to create our Dome Phalanx Finger (DPF) topology (see Figure 2a). The finger derives its behavior from the mechanical response of the dome-shaped shell elements and the interaction between each segment. As the units invert, the top section of the finger expands, while the limiting layer (see Figure S1d) retains its original length. This results in global curvature and different final stable shapes, depending on the number of inverted dome units (see Figure S2b).
These dome-shaped units program different mechanical responses and stable shapes into the system, effectively discretizing the infinite-dimensional deformation space into a manageable number of kinematic configurations dependent on the dome unit geometry, each attainable via well-defined, open-loop inputs. The mechanical response and stable configurations of the DPF can be tuned by adjusting the dome height per unit ($H_{i}$), dome thickness ($t$), number of segments, dome unit size (UC), pneumatic chamber thickness ($t_{\text{ch}}$), limiting layer thickness ($t_{\text{lim}}$), dome unit length ($U^{i}_{\text{L}}$), air chamber dimensions ($W_{\text{ch}}$ and $t_{\text{mid}}$), and the spacing between adjacent cells per unit ($U^{i}_{\text{sep}}$) (see Figure 2a and Figure S1a for reference). The geometric parameters of the dome units $H$ and $t$ dictate the mechanical response, determining whether it is monostable (Figure S1c(i)), metastable (Figure S1c(ii) and Movie 1), or bistable (Figure S1c(iii)). Furthermore, the dynamic response of the DPF can be controlled by the material properties and additional geometric parameters ($t_{\text{lim}}$, $t_{\text{ch}}$, $t_{\text{mid}}$), which significantly influence global curvature, system overshoot, and steady-state error (see Figure S3).
Given the broad range of geometrical configurations, each finger can be geometrically tuned to exhibit different mechanical responses: a bistable response, where the final shape is retained after dome inversion [23, 17, 46] (Figure 2b $H_{i}$ = 4 mm and $H_{i}$ = 5 mm). A metastable response, where each unit undergoes snap-through after which its shape is retained for a time $t=\tau$, before returning to its stress-free state as a consequence of the combined effects of the unit’s viscoelastic material behavior and dome geometry [36, 47, 48] (Figure 2b, $H_{i}$ = 3 mm and Movie 1). Lastly, a monostable response, where the material behavior dominates over the geometric response and the finger responds similarly to a PneuNet bending actuator. As a result of the encoded mechanical responses, the actuation pressure required to invert the dome units is decoupled from the final shape, which is instead entirely determined by geometric parameters. Figure 2 d presents experimental measurements of the DPF tip displacement magnitude over time for various actuation pressures (dashed lines), demonstrating that the same final position is achieved despite different actuation magnitudes (Input pressure). Moreover, we leverage the material’s viscous response and structural properties to reset the system, whereby the designed unit metastability (dependent on dome height, material properties) and the duration of the applied load control the reset dynamics (see Figure 2c). The resulting freedom to design the DPF’s dynamics and set points allows for pre-programming a diverse set of tasks, sequences, and positions.
Figure 2: Dome Phalanx Finger (DPF) with different encoded mechanical responses and behaviors. a) 5-segment DPF geometry parameters ($i=5$). b) Effect of dome height on finger response (Every segment with the same height). $H_{i}$ = 5 mm and $H_{i}$ = 4 mm show a bistable behavior, and $H_{i}$ = 3 mm shows a pseudo-bistable (metastable) response (self-resetting system). c) Bistable finger ($H_{i}=5$ mm) Tip displacement and pressure over time. d) Metastable finger ($H_{i}=3$ mm) Tip displacement and pressure over time.
e) Time lapse of two gripper architectures showing both bistable and pseudo-bistable responses.
To showcase these unique capabilities, we developed a two DPF gripper architecture. By changing the dome height, we can program a gripper composed of two fingers to pick and hold an object after actuation (Figure 2c, $H_{i}$ = 4 mm) or pick an object and release it as the dome units reset (Figure 2c, $H_{i}$ = 3 mm), the latter of which is ideal for embodied pick-and-place task typically performed by conventional soft grippers. The number of stable states that can be encoded using the DPF morphology equals that of the number of constitutive units where each unit is tunable independently. Consequently, different unit geometries can be combined to achieve multiple stable states enabling a diversity of behaviors within the same overall geometry (Bistable + Metastable DPF). This allows for different energy minima and time responses to be embodied within the same finger, which can be accessed by a single pressure input and serve as discrete and functional robot configurations. To effectively design, combine, and predict the finger’s behavior based on its geometry, a simple yet reliable modeling framework is necessary. Given the broad design space and the computational cost of using FE models for our finger topology, we propose an energy-based spring model that captures the complex mechanical behaviors needed to design multistable soft robots with desired functional characteristics.
2.2 Model derivation and mechanics-informed parameter determination
Establishing an analytical model to capture the global behavior of the DPF requires consideration of both the unit cells’ and the global geometric parameters. By capturing the contribution of each constitutive unit of the DPF and their interactions, we can construct an energy landscape with a simplified model, where each of the minima corresponds to a stable state and a final position of the system. In this way, the state of the system ($x_{i}$) can be extracted from the resulting strain energy potential via a minimization process, whereby the DPF’s geometrical and stiffness characteristics dictates the programmed stable shapes. The DPF geometry is represented by a lattice array comprising nonlinear, linear, and torsional springs for the static response and point masses and dash-pots for the dynamic response (Figure 3a). The energy for the springs used in the lattice can be written as follows [49]:
Nonlinear: $\text{{Nonlinear:} }E_{NL}(x_{ij})=\frac{1}{2}k_{b}x_{ij}^{2}\left(1+\left((1-\alpha)\left(\left(\frac{x_{ij}}{d}\right)^{2}-2\frac{x_{ij}}{d}\right)\right)\right)$ (1)
Linear: $\text{{Linear:} }E_{L}(x_{ij})=\frac{1}{2}k_{l}\cdot x_{ij}^{2}$ (2)
Torsional: $\text{{Torsional:} }E_{T}(\vartheta)=\frac{1}{2}k_{\vartheta}\cdot\vartheta^{2}$ (3)
$E_{tot}(x_{ij})=\sum_{i=1}^{n}E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij}).$ (4)
where $x_{ij}=x_{i}-x_{j}$, $\vartheta$ is the angle between the rigid link and the limiting layer (see Figure S5a and S7c) and $k_{b}$, $k_{l}$, $k_{\vartheta}$ are stiffness parameters that depend on the DPF geometry. The individual stiffness and spring connectivity allow us to capture the limiting layer and dome inversion effects. We use a nonlinear spring to represent different dome mechanics (Figure S1c), which are described by the parameters $k_{b}$, $\alpha$, and $d$. Linear and rotational springs are utilized to capture the limiting layer of the finger by combining stretching and bending energies (see Figure 3a). The connection between the limiting layer and the nonlinear springs is modeled as an infinitely rigid connection, given that the strain is negligible compared to the rest of the lattice elements. By minimizing the sum of all contributions to the system (Equation 4), the static problem is solved, and all possible finger stable configurations are predicted (see Figure 3b).
A primary challenge of spring-lattice models is to accurately map the geometric parameters of the system to its stiffness constants. To address this challenge, we integrate FE simulations (see section A.2) with Recursive Feature Elimination (RFE; see section A.3[50] and lasso regression to derive expressions for $k_{b}$, $\alpha$, and $d$—the nonlinear spring parameters in Equation 1—as functions of dome height ($H$), dome thickness ($t$), dome curvature ($R$) and Young Modulus ($E$) of a generic constitutive unit (see Figure S1). We considered different dimensionless relations ($\pi_{i}$) based on the most relevant geometric interactions affecting the dome’s energy [23]. These mechanics-informed interactions enhance the generality of the analysis by linking the mechanics to the unit’s geometry. For our model, we assume all spring constants can be expressed as functions of the shell’s load-carrying capacity ($\pi_{1}=\frac{t}{H}$), curvature-to-thickness ratio ($\pi_{2}=\frac{t}{R}$), and dome shallowness ($\pi_{3}=\frac{H}{R}$). Consequently, each parameter can be formulated as:
[…]
where $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ is a matrix containing all possible candidates and interactions of the non-dimensional relations, as shown in Equation S.14. Using RFE, the relevance of each feature can be automatically determined by iteratively removing one feature at a time and observing the model’s coefficient of determination ($r^{2}$). This process yields the weights $\xi_{\alpha}$, $\xi_{k}$, and $\xi_{d}$ with fewer features while maintaining high accuracy. Notably, only the unit cell is used to determine the parameter features, and as a consequence, the DPF’s mechanical behavior is derived. A wide range of Young’s modulus ($E$), dome height-to-dome base ratio $\left(\frac{H}{r_{b}}\right)$, dome base radius ($r_{b}$), and thickness ($t$) are simulated to capture the entire design space of the DPF, including various soft materials ($5MPa\geq E\leq 40$ MPa). As a result, we derive a general equation in terms of the three non-dimensional parameters (Equation S.16, Equation S.17, and Equation S.18, see Section A.3.3 for details). Using these derived expressions and a gradient-based optimization algorithm (see Section A.4 for more information), the final shape of the finger can be predicted by minimizing the system’s energy, which is now expressed as a function of the geometric parameters of the structure.
Figure 3: An energy-based model for static and dynamic analysis of multistable soft robots based on the DPF topology. a) Spring lattice model with linear, rotational, and nonlinear springs. b) Static stable states predicted by the lattice model for $H_{1}$ = $H_{2}$ = 4 mm, $H_{3}$ = $H_{4}$ = 4.5 mm, and $H_{5}$ = 5 mm. c) Dynamic response for the soft robot with four bistable units and two metastable units ($H_{1}=H_{2}=H_{3}=H_{4}=5$ mm, and $H_{5}=H_{6}=3$ mm). The model captures the snap-through of each bistable unit and the snap-back of the metastable units (See Movie 2).
A comparison between FE simulations and the proposed model for different stable states of the finger (see Figure 3b) and different design cases (see Figure S10) shows good agreement with the energy-based lattice model, while reducing the computational time by more than $10^{-3}\times$ (see Table S2). The model is further extended to capture the structure’s dynamic response, allowing us to observe the material’s viscoelastic behavior and analyze metastable cases that energy minimization alone cannot predict. The dynamic response for different unit cell geometries is shown in Figure S7b, and for different five-segmented fingers in Figure S7d, where the dynamic model predicts the same final state of the system after reaching a steady state. To explore the combined behavior of different dome geometries, we modeled a five-segment DPF topology with bistable and metastable units (see Figure 3c). We stacked four bistable units ($H_{1}=H_{2}=H_{3}=H_{4}=5$ mm) and two metastable units with the same dome height ($H_{5}=H_{6}=3$ mm). This configuration allows us to achieve one unique set point: a stable state and a time-dependent state that holds its shape for a given predictable time $\tau$ (see Figure 3c). The resulting dynamic response features five different snap-through events and two snap-backs (dome-unit resets), adequately representing the complex dynamics displayed by the DPF (Figure 3c).
2.3 Position and stiffness inverse design
Based on the model’s performance and its agreement with the FE simulations, we pose an inverse problem to co-design different DPF topologies capable of reaching a desired pre-programmed final position to achieve embodied position control. Since the stable positions of the system are already determined by the stable states, the optimization algorithm only searches within the fully actuated configurations, making it computationally efficient. We utilized Bayesian optimization [51] to iterate over various numbers of segments, dome heights for each of the segments ($H_{i}$), thickness of the dome ($t$), unit separation for each of the segments ($\text{U}^{i}_{\text{sep}}$), and unit length for each of the segments ($\text{U}^{i}_{\text{L}}$) (see Figure 2a for parameter reference). Two different objective functions are utilized: the first based on position and targeting a specific tip coordinate (Position Design), and the second one targeting a specific tip coordinate while maximizing the finger’s stiffness (Position + Stiffness Design; see section A.6).
Position Design: Given a target tip coordinate [x,y] ($\text{Target}_{xy}$), the inverse problem objective function can be written as:
$\min_{H_{i},t,\text{U}^{i}_{L},\text{U}^{i}_{\text{sep}}}\left(\text{Target}_{xy}-\text{Tip}_{\text{dis}}\right)^{2}$ (6)
[…]
$\hskip 14.22636pt\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N$
$\hskip 14.22636pt\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N$
where the $\text{Tip}_{\text{dis}}$ represents the tip point (see Tip in Figure 2a), which coincides with the top node of the finger’s last unit. A domain constraint is imposed to guarantee that every dome unit is higher than the one in front ($H_{i+1}\leq H_{i}$), which in certain cases can lead to up to N possible stable states (N been the number of segments). Moreover, we impose constraints for all unit separations and unit lengths to be equal for simplicity and design domain to guarantee bistability on all units, based on the phase diagram shown in Figure S1b. However, different restrictions can be imposed depending on the DPF design purpose, and parameters can be individually tuned to achieve a large variety of positions and stable states (see Movie 3).
Our algorithm was tested by defining five different tip coordinate targets using two materials with distinct properties (NinjaFlex 85A and Cheetah 95A) optimized to achieve the desired positions. For each case, the target coordinates in $x$ and $y$ were provided to the algorithm (Target Position in Figure 4a), and the optimization was performed by incrementally adding one unit per run and calculating the objective function (Equation 6). To reduce the computational cost, a maximum of 8 units was specified (Figure 4b), although additional units can be included if the objective requires a larger finger geometry. The geometry with the minimum objective function value (Obj in Tables S3 and S4) was selected and plotted to verify whether the desired position was achieved (see Figure 4a).
Geometric parameters, coordinate targets, and objective function values are reported in Table S3 and S4. The final stable state and its comparison with each target can be observed in Figure S11. To validate our optimization algorithm, we 3D printed the five different geometries shown in Table S9 using both tested materials (NinjaFlex 85A and Cheetah 95A) and measured the final tip position after commanded dome inversion. An average error of 7.3% was measured across all samples, showing good agreement with the optimization algorithm and the lattice model.
The topologies generated by this methodology exhibit deterministic behavior despite the use of polymeric materials, meaning that kinematic states are consistently attained as long as the gripper’s geometric parameters remain unchanged (see Figure 4c and Movie 4). Specifically, we cycle each gripper topology between the initial and fully activated states, verifying the gripper’s ability to achieve the same aperture and positioning across multiple cycles (Tests 1-3 in Figure 4c). Furthermore, we assessed material and time degradation effects by performing 300 cycles over three weeks (see Figure S17), which demonstrates that even after numerous activations and an extended period, the gripper’s performance remains within 5% of the initial design behavior. This characteristic enables robust robot design, as the bistable elements provide intrinsic kinematic control dictated by the gripper’s mechanical response. Notice that the pressure for activating the units is not controlled using sensors and closed-loop control, illustrating the embodied self-regulation from the designed multistability. Consequently, the position control inherent to the designed multistable structure serves as control set points that can be embodied in the robot’s morphology and achieved with a single, open-loop controlled actuator.
Figure 4: DPF inverse design and experimental validation. (a) Predicted results for Target 1 and Target 3 from the optimization model (Position Design). (b) Objective function values as a function of the number of segments, with the minimum value identified for each target position. (c) The gripper achieves multiple stable kinematical states predicted by the energy-based model, which are successfully transferred to 3D-printed prototypes. (d) Stiffness determination of the DPF using a follower force approach, with Normalized strain energy vs Displacement curves shown for Objectives 1, 3, and 5. (e) Comparison of the proposed model and experimental results for position-only (Pos) optimization and combined position + stiffness optimization (Stf).
Position + stiffness design: The finger’s stiffness is calculated by using the dynamic model described in section A.4, by perturbing the structure from a programmed stable configuration (i.e., set point). We determine the system’s total internal force by solving the dynamic equation:
$[M]\ddot{x_{i}}+F_{d}^{int}(x_{ij},\dot{x_{i}},\dot{x_{j}})+F_{d}^{iso}(\dot{x_{i}},\dot{x_{j}})+F_{\text{in}}=F_{ext}(x_{i,j},t)$ (7)
$F_{\text{in}}=\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)$
where $F_{\text{ext}}(x_{i,j},t)$ is a follower force applied at the tip of the DPF (see Figure 4) to perturb the DPF from its equilibrium position. The gradient of the internal forces in the structure, $F_{\text{in}}$, is obtained through symbolic differentiation of the expression described in Section 2.2. The force is applied at a constant load rate of 5 mm/min, ensuring a quasi-static regime. The stiffness of the DPF in the desired state is calculated by determining the internal force ($F_{\text{in}}$) and the corresponding tip displacement over time (see Figure 4). A linear fit is then applied to the resulting data. Using this methodology, the DPF’s geometry can be optimized to achieve a specific target position with maximum stiffness in the vicinity of that stable state (see Figure 4e). Equation 6 can thus be reformulated as a weighted optimization problem expressed as:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s}}w_{1}\left(\text{Target}_{xy}-\text{Tip}_{\text{ dis}}\right)^{2}+w_{2}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}$ (8)
[…]
$\hskip 14.22636pt\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N$
$\hskip 14.22636pt\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N$
where $w_{1}$ and $w_{2}$ are the weights assigned to each objective, and $K_{\text{dis}}$ is the stiffness obtained from position inverse design (Equation 6). The number of segments (N) is kept consistent with the results obtained from the previous optimization (Equation 6), meaning that $H_{i}$, $\text{Unit}^{i}_{L}$, $t_{\text{lim}}$ and $\text{Unit}^{i}_{s}$ are considered to maximize stiffness. The results of the stiffness and position optimization are shown in Table S6 for Cheetah 95A and in Table S5 for Ninjaflex 85A. Moreover, we experimentally test Target 5 geometries produced by both optimization procedures (see Figure S16 and Movie 5). By following this approach, we had increased the DPF initial stiffness on average 1.94 times for all cases, achieving maximum stiffnesses of 0.55 N/mm, 0.3 N/mm, and 0.15 N/mm for cases 2, 4, and 5 respectably (see Figure 4e for case 5), which consequently enhances the achievable grasping force.
2.4 Dome Phalanx Robot (DPR) performance
To explore the capabilities of our design approach and the advantages of leveraging multistability in soft robotics, we fabricated a soft robot (Dome Phalanx Robot - DPR) by combining four DPFs ($A_{1}$ to $A_{4}$), as shown in Figure 5a. We assembled the four DPFs using 3D printing a rigid base that prevents interaction between the fingers, imposes appropriate boundary conditions our model considers, and couples our robot to a UR3 robotic arm (see Movie 6). This robot can be assembled by either combining fingers with the same topology ($A_{1}=A_{2}=A_{3}=A_{4}$), resulting in uniform global functionality, or by using two sets of actuators ($A_{1}=A_{3}$ and $A_{2}=A_{4}$), which provides multiple functionalities within the same robot body.
We demonstrate different operation modes using the proposed design methodology and FDM 3D printing to create multiple robot topologies. Specifically, we focus on three operation modes: 1) A multistable robot with at least three set points/stable states (see Figure 5a); 2) A configuration with two finger pairs that can exert different grasping forces but reach the same final position (see Figure 5b); and 3) A combination of bistable and metastable dome geometries (see Figure 2c and d), where the DPR’s constitutive material viscoelastic response of the material allows us to embody a pick and place robotic behavior into the robot geometry (see Figure 5c and d and Movie 7). For each of these operation modes, different constraints (detailed below) are imposed on the model to ensure that the structure is accurately designed for each specific task.
Figure 5: Dome Phalanx Robot (DPR) assembly and modes: Fingers 1 to 4 ($A_{1}$ $A_{4}$) can be changed for different operation modes and mechanical behaviors. a) Multistable robot with four different set points: (i) Stable state 1 with 0 inverted units, (ii) Stable state 2 with two inverted units, (iii) Stable state 3 with four inverted units, and (iv) Stable state 4 with six inverted units. b) $A_{1}$ and $A_{3}$ are optimized for position and stiffness, while $A_{2}$ and $A_{4}$ are optimized for position. $A_{2}$ and $A_{4}$ exert $F_{Dis}$, and $A_{1}$ and $A_{3}$ exert $F_{\text{stt}}$ ($F_{\text{stt}}>F_{\text{Dis}}$). c) Bistable + Metastable finger behavior: Strain energy and pressure vs. time response. d) Snapshots of embodied pick-and-place tasks: A single pressure input is given for the robot to pick the object (Commanded response) and then release it depending on the viscoelastic material response (Programmed response)(see Movie 7).
1) Multistable soft robot: We further leverage the characteristics of our multistable DPF by programming different set points for the same robot morphology (i.e., DPFs geometries). Using the optimization algorithm, we can program three stable states by minimizing for three discrete configurations (2, 4, and 6 active units). This design approach can be implemented by modifying the objective function as follows:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s}}w_{j}\left(\text{Target j}_{xy}-\text{Tip}_{\text{dis j}}\right)^{2}+w_{M+1}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}$ (9)
[…]
$\hskip 14.22636pt\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N$
$\hskip 14.22636pt\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N$
where $M$ is the number of targeted stable states, and $j$ is the stable state number ($0<j\leq M$). By doing this, we can optimize for three target positions and maximum stiffness in the fully activated state (six inverted units). For this design, we impose the number of set points, the target tip position for each set point, and the number of units. It is worth noting that for this case, there is a different constraint for the dome height (i.e., $H_{i+1}>H_{i}$), which allows each of the states to be useful for grasping objects while being accessible in a sequential manner. While the model could be expanded to modify the unit cell’s dimensions further and determine the optimal number of units, for simplicity, we fixed the number of units at six to ensure at least two active domes per stable state. We tested our system by performing an inverse design with three different target positions (see section A.6.3 for reference), using the same methodology described previously (section 2.3). An average error of 12% was found between the desired positions and the three final stable states, as shown in Figure 5a ((i) - (iv)).
This methodology demonstrates the versatility of our DPF and provides an alternative approach to encoding multiple set points within the finger’s body by modulating its unit cell geometry. The local geometric modifications create more accessible stable states, which can be controlled using a single, open-loop driven pressure inlet per DPF. Since the systems’s curvature and stiffness are functions of the number of active units (see Figure S15b), the grasping force and carrying load capacity of the DPR also become functions of the active domes. Consequently, the robot can also be programmed with different set points for force (see Movie 6). Moreover, to showcase the gripper’s versatility to grasp various objects and its carrying capacity, we performed grasping of different objects on different stable states of the DFR (see Figure S18 and Movie 6). Notably, we observed that the restriction of the object can reset some of the dome units, which makes the DFP adapt to the object shape and perform a better grip depending on the object size. A maximum load-to-weight ratio of 10.25 is determined by loading the robot until it can not withstand the weight (see Movie 6).
2) Force-adaptable robot: For the second operation mode, we utilized the results shown in section 2.3, where two different objective functions were considered to create two different finger topologies that achieve the same tip position but exert different grasping forces. Using the data from Table S3 and Table S5, we designed a robot arrangement where two fingers ($A_{1}$ and $A_{3}$) produce more grasping force than the others ($A_{2}$ and $A_{4}$). We tested the grasping force using a compliant mechanism composed of two bars that compress as the grasping force increases (see Figure 5b). We characterized the mechanism’s force vs. aperture to measure the robot’s grasping force (see Movie 5). Our experiments show that the two different objective functions result in a force increase of 1.8 times ($F_{{\text{Dis}}}$ and $F_{{\text{stt}}}$) while reaching the same position in their fully activated stable state.
3) Embodied soft robotic tasks: In the third operation mode, we leverage our ability to program our system’s dynamics. By leveraging our dome structures’ viscoelastic response and geometry, we can pre-program the duration for which the structure remains activated before returning to its original state. As demonstrated in section 2.1, we can modulate the geometry to create a fully metastable finger (see Figure 2d and e). However, the same finger can exhibit bistable and metastable behaviors by combining bistable and metastable dome units. This approach allows for encoding a desired set of points into the finger and operate in the vicinity of that point by actuating the metastable unit cells (See Movie 2). Combining the meta- and bi-stability of the units enables us to design task planning operations (such as pick-and-place; see Movie 7), where the robot can pick an object by activating all units, move to a specific position, and release the object without additional interventions and with just one pressure input (i.e., eliminating the need for a release command; see Figure 5d). We tested this operation mode by designing a six-segment finger with four bistable and two metastable units. The set point with four active units is designed with $H_{1}=4$ mm, $H_{2}=4.5$ mm,$H_{3}=5$, and $H_{4}=5$ mm, while the metastable units are designed to maximize the resetting time for two different loading times (section A.6.4). We approximate the loading profile, by measuring the pressure experimentally and approximating it as shown in Figure S13. The obtained results show how an object can be held while all units are inverted, even when the system releases pressure (Figure 5c). However, the object is automatically released from the gripper after the metastable units reset, as desired from the programmed dynamics. Using this approach, we can program a simple pick-and-place operation by activating all domes or perform a simple pinch operation as demonstrated in Figure 5d and Movie 7. Further flexibility in the attainable behavior of our robots can be obtained by including the actuation loading time, which results in the same DPR morphology adapting its response to can pick an object and release it at two different times as demonstrated in Movie 7.
3 Conclusion
In this study, we introduced a modeling framework and inverse design methodology for soft robotics, discretizing their continuous response using multistable structures. We propose a robotic architecture, The Dome Phalanx Finger, that leverages mechanical multistability to discretize the robots’ configuration space into desired kinematics, which serves as control set points. Exploiting these characteristics, we embodied diverse static and dynamic responses providing an alternative open-loop control strategy for realizing various robotic tasks normally requiring closed-loop approaches guided solely by the designed geometry and material response. Implementing an energy-based modeling approach was central to our methodology, enabling us to accurately predict and program the DPF’s behavior. Through this approach, we established a robust framework for modeling the finger’s static and dynamic responses, with a mechanics-informed parameter discovery procedure that yields high-fidelity results consistent with experimental data. Significantly, this methodology extends beyond the DPF geometry and its lattice representation, as the combination of mechanics knowledge (in this case from shell theory), data, and recursive feature elimination can be further exploited for a wide variety of geometries and reduced-order models. Utilizing the predictive capabilities of our model, we successfully implement the inverse co-design of the DPFs’ morphology and controlled functionality yielding desired target positions, accessible states, and programmable time responses based on the structure geometry and material viscoelasticity. Our results show how these characteristics can be designed to allow for the modulation of grasping force and unit reset without the need for closed-loop control. Our findings highlight the potential gains in performance when morphology and functionality are concurrently optimized (i.e., co-designed), a result only feasible via the simplified dynamics and modeling afforded by our approach. The implications of our results are far-reaching, suggesting a promising path toward developing robust soft robots. By encoding the control within the morphology and response of the system, we embody a form of mechanical intelligence offering a reduction in control complexity and forgoing the need for data-hungry models and expensive electronics. More importantly, this approach is not only beneficial for systems controlled in open loop. As the robot behavior is transformed into a finite, tractable set of positions or set points, the implementation of closed-loop control strategies is simplified by the tracking of a finite number of states instead of the infinite number inherent to continuous systems. Combining our morphological control strategy with conventional sensors can create a streamlined closed-loop control, where sensors can be utilized to verify if the robot is in a specific state, thus reducing the need for constant feedback from the system. The resulting reduction in system complexity paves the way for more efficient, cost-accessible, sustainable, and adaptable soft robotics applications.
4 Experimental Section
Unit cell and DPF Fabrication– The unit cells are 3D printed using fused deposition modeling (FDM) on a Raise3D Pro2 printer, utilizing Ninjatek Ninjaflex TPU filament [52] and Ninjatek Cheetha TPU [53]. Each finger is post-processed to correct manufacturing errors and seal the air chambers. For the DPG, we printed a PLA base that secures each of the four fingers (see Movie 6 for assembly).
Stiffness Measurement tests– The actuators were loaded onto a specially designed test rig and tested using a universal testing machine (3345 Instron) equipped with a 100N load cell (see Figure S16a and Movie 5). The test rig constrained all degrees of freedom at the finger’s base, with six different positions equally spaced along 90° to ensure that the initial load was always perpendicular to the tested geometry. We measured the force-displacement response five times per sample at a constant velocity of 10 mm/s. Raw data was evaluated and plotted using Python 3.12.
Finite element analysis– The unit cells are modeled in Abaqus using linear elastic material properties and a combination of S3R and S4R shell elements. A dynamic implicit quasi-static approach is utilized to capture the structure’s instabilities. Geometric nonlinear analysis (Nlgeom) is used, and snap-through is triggered using a pressure load on the dome while the edges are pinned. The unit cell is initially modeled in the stress-free state. Simulations were run for different material thicknesses ($t$) and dome heights ($H$). The full geometry of the DPF is also modeled in Abaqus using linear elastic material properties and C3D10 3D elements. Snap-through is triggered using a pressure load on every inner wall air chamber. As discussed in the analysis, a fixed boundary condition is applied to the first unit. After dome inversion, different relaxing steps are used to achieve the final stable state (see section A.2 for details).
Robot actuation– Each DPF is actuated using a pressure input controlled by a ball valve. The results displayed in Movies 2 - 3 are produced by applying pressure as slowly as possible to observe the multistable phenomenon.
Pressure and displacement measurement– The dynamic behavior of the DPF is captured in detail using a Photron Fastcam Mini UX100 high-speed camera and the data is processed using Tracker Video Analysis and Modeling Tool. Input pressure is measured using a HONEYWELL ABPDANN010BG2A3 pressure sensor coupled with an Arduino Uno (see section A.7.1 for more details). Each experiment is performed 5 times and the average values are reported.
Acknowledgements
The authors gratefully acknowledge the support from the DARPA Nature as Computer (NAC) program under the guidance of Dr. J. Zhou and NSF-CAREER award No. 1944597.
Appendix A Supplementary Material
A.1 Geometry
A.1.1 Unit Cell
The system’s bistable constitutive units are composed of a dome-shaped structure encapsulated by a square chamber that allows for pneumatic actuation and reset by applying positive and negative pressure, respectably (see Figure S1a). Each dome unit can be geometrically tuned to exhibit a bistable behavior (Figure S1c (iii)) [23], pseudo-bistable (metastable - Figure S1c (ii)) behavior where the is a snap-through instability, but the unit returns to its zero energy state [47, 48], and a monostable behavior (Figure S1c (i)). These mechanical responses can be adjusted by modifying the height $H$ and thickness $t$ of the unit, as shown in Figure S1b.
Figure S1: Dome unit geometry and mechanics. a) Unit cell geometry and parameters (Dome unit + air pressure chamber). b) Stability map dependent on dome thickness ($t$) and dome height ($H$) for DPF unit cell. c) Monostable (i), Metastable (ii), and Bistable (iii) dome unit behavior. d) Dome Phalanx finger geometry with two possible, stable states (Initial and fully actuated states).
A.1.2 Dome Phalanx Finger (DPF)
The DPF is multistable soft system with sequential dome units (Figure S1d). The finger derives its multistable behavior from domed-shaped shell elements (Figure S1a), and it can be geometrically tuned to reach and retain different final shapes after dome inversion. The DPF performance and stable kinematic configurations can be tuned by changing the dome height ($H$), Unit Cell size (UC), strain limiting thickness ($t_{\text{lim}}$), air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$), spacing between adjacent cells ($\text{U}^{i}_{\text{sep}}$), unit cell length ($\text{U}^{i}_{\text{L}}$), chamber thickness ($t_{ch}$) and dome thickness. Once fully inverted, contact between adjacent units and the strain-limiting layer induces the system’s global curvature. The domes on each finger segment support programmable deflections as the domes’ final positions dictate the global kinematic configuration (see Figure S2b). The contact between the dome tip and the adjacent unit’s chamber provides an additional interaction, contributing to the final curvature after all units are activated and the pressure is removed. The strain energy stored in the gripper topology in its activated states is the summation of the individual contributions from the inverted domes (see Figure S2c). The magnitude of the interaction between neighboring domes is also affected by dome height ($H$) and chamber thickness ($t_{ch}$). Together, dome height and chamber thickness determine the degree of interaction between neighboring segments after dome inversion and yield the system’s global curvature (Figure S2b) and the final tip position of the system (Figure S2c).
Figure S2: Dome Phalanx Finger geometry and behavior. a) Geometrical parameters that determine the final position and strain energy of the DPF. b) Different stable states of the DPF are encoded by tunning the dome height of each unit cell. c) Strain energy (i) and Tip displacement (ii) as a function of the number of active units. Three different sets of points are encoded into the DPF.
Consequently, various dynamic behaviors can be achieved by adjusting the geometrical parameters, thereby enhancing tunability and control. Specifically, the dome height ($H$) and thickness ($t$), along with the limiting layer thickness ($t_{\text{lim}}$), influence the final stable position or set point. Meanwhile, the chamber thickness ($t_{ch}$) and the air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$) govern the overshoot required to reach a target position (see Figure S3). This tunability enables the gripper’s response to be tailored by simply modifying the DPF’s geometric characteristics.
Figure S3: Geometric parameter effect on the time response of DPF. Setpoint, overshoot, and dynamic response can be tuned by combining different parameters. a) Elastic Modulus (E), b) Dome unit Height (H), c) Chamber thickness $t_{ch}$, d) Limiting layer thickness $t_{\text{lim}}$, e) Dome unit thickness ($t$), f) Channel with $W_{\text{ch}}=k_{\text{mid}}UC$
A.2 Finite Element Simulations
Finite element simulations are performed using Abaqus and Python scripting to iterate over different geometrical configurations. Given the dominance of the geometric phenomena in the system, all simulations are done using linear elastic material properties.
A.2.1 Unit Cell Simulation
The unit cells are modeled using S3R and S4R shell elements. The mesh is structured to capture the dome symmetry and its appropriate post-buckling behavior. A dynamic implicit quasi-static approach is utilized to capture the structure’s instabilities. Geometric nonlinear analysis (Nlgeom) is used, and snap-through is triggered using a displacement control method on the dome while the edges are pinned. The unit cell is initially modeled in the stress-free state. Simulations were run for different material thicknesses ($t$), dome heights ($H$), dome radius ($R_{b}$), and Elastic Modulus ($E$). Strain energy vs dome tip displacement is extracted for every step of the simulation (see Figure S4a) which are used to tune the lattice model constant ($k_{b}$ and $\alpha$ in Equation 1).
A.2.2 Dome Unit interaction
The interaction between dome units is modeled by simulating two consecutive units (see Figure S4b). This approach allows us to assess the influence of one unit on the final position of the adjacent dome tip. The interaction is captured using general contact between the surfaces while varying key parameters such as material thickness ($t$), dome height ($H$), dome radius ($r_{b}$), and elastic modulus (E). Tip displacement is recorded over time, and the final displacement values, as a function of the test parameters, are used to calibrate the parameter $d$ (Equation 1) in the lattice model.
A.2.3 DPF 3D Simulation
The complete geometry of the DPF is modeled using C3D10 3D elements to capture the detailed deformation behavior. Snap-through is initiated by applying a pressure load to each inner wall of the air chambers. 3D simulations are chosen to capture the dynamic behavior under internal pressure loading. A fixed boundary condition is applied to the first unit to prevent rigid body motion. After dome inversion, a series of relaxation steps are performed to allow the structure to reach its final stable state (see Figure S4 c). To validate the lattice model, simulations are conducted for various geometric parameters, including the number of units ($N_{\text{Units}}$), material thickness ($t$), chamber wall thickness ($t_{\text{ch}}$), elastic modulus (E), dome height ($H$), and base radius ($r_{b}$). The final deformed shape is extracted for comparison with experimental results. The entire DPF geometry and simulation workflow is implemented using Python scripting in Abaqus, ensuring efficient parameter exploration and model validation.
Figure S4: Finite Element (FE) simulations for model parameter tuning and validation. (a) Dome unit inversion sequence used to determine $k_{b}$ and $\alpha$. (b) Interaction between dome units to determine $d$. (c) 3D simulation for validating the model’s predictions against FE data.
A.3 Energy-based model and constant derivation
A.3.1 Spring Lattice array
Predicting the behavior of multistable DPFs requires consideration of both the unit cells’ and the global geometric parameters. By characterizing the contribution of each DPF subcomponent and their interactions, an energy landscape can be built where each minimum corresponds to a stable state of the system. The resulting strain from the energy minimization process dictates the programmed stable shapes with their geometrical and stiffness characteristics. We represent the DPR as a lattice comprising nonlinear [49], linear, and torsional springs (Figure S5a). The springs’ stiffness and connectivity allow us to map local extensions and rotations to defined energy contributions. We use nonlinear springs featuring a ground (unstressed) state and an inverted (stressed) to capture the influence of the bistable domes on the system. To appropriately represent the bistable dome behavior, we position the nonlinear springs so that the path of extension coincides with the dome’s tip position between the ground and inverted states. Linear springs capture connections between the nonlinear units modeled as struts. In the DPF’s case, the strain these segments experience is negligible, behaving as a strain-limiting layer. Consequently, we model these connections as nearly rigid. While the axial strains experienced by this strain-limiting layer are negligible, the layer is sufficiently thick to display bending resistance. We capture the influence of this bending stiffness by introducing torsional springs at the nodes coincident with the strain-limiting layer. The angular displacement experienced by the torsional springs is dictated by the change in the angle $\vartheta$ formed by the segments (see Figure S5a).We define $k_{l}=\frac{\sqrt{3}}{2}E\frac{t_{\text{lim}}UC}{U_{L}+t_{\text{ch}}+U_{\text{sep}}}$ and $k_{t}=\frac{2}{\sqrt{3}}\frac{E(t_{lim}+t)}{12(1-\nu^{2})}$ Where $E$ is the young modulus, $\nu$ is the Poisson ratio, and $t_{lim}$ is the thickness of the limiting layer shown in Figure S2a.
Minimizing the system’s total energy, $E_{tot}=\sum_{i=1}^{n}E_{L}+E_{NL}+E_{T}$ where $n$ is the number of units, constrains the space of possible interactions for the lattice elements into a discrete set of stable states. The rich configuration space of our DPF requires establishing a rational method for providing the initial guess and initializing the minimization process. To this end, we implement a geometric base model to generate initial guesses for the optimization algorithm so that the obtained states are in the neighborhood of physically feasible configurations with improved computational time. The geometric model for the dome phalanx finger assumes a fixed extension of the nonlinear spring and pure rotation from segment to segment. Here, the activated states of the finger (i.e., with units with inverted domes) are determined by extending the length of the linkage corresponding to the bistable dome structure such that it reflects the distance along the central axis between the tip of the dome to the base of the dome chamber. The now extended segment yields an angle $\vartheta$ between the present and initial position of the front segment (see Figure S5 a). The model uses this $\vartheta$ to calculate the rotation of the subsequent segment, and the process is repeated until the number of segments is exhausted. In this model, the displacement of the overall fingertip can be calculated as a summation of sines and cosines. This purely geometric model does not account for the internal force balancing characteristic of each stable state, which ultimately leads to errors in the configuration predictions. Nevertheless, predictions of a purely geometric model are a powerful tool in finding solutions within the hyperdimensional energy landscape of the spring lattice when we use the outputs of the geometric model to form the initial guess in our search for local extrema. The stable configurations that result from balancing the force and energy contribution of coupled springs differ from the activated configurations determined by the geometric model, but the discrepancy is one easily overcome by the energy minimization process.
A.3.2 Material viscoelastic response
Material time response is captured using a variation of the Kelvin-Voigt viscoelastic constitutive model[54]. The model consists of two components in parallel: a spring (representing the elastic behavior) and a dashpot (representing the viscous behavior) see Figure 3a.
Nonlinear Spring (Elastic Element): The spring represents the material’s dome behavior, as explained in the previous section. This spring follows Equation 1.
Dashpot (Viscous Element): The dashpot represents the material’s viscous behavior, meaning it resists deformation by dissipating energy as heat. This behavior is modeled by considering two different damping forces, isotropic damping, and internal damping [55]. The isotropic damping for numerical stability on a node $i$ can be written as:
$F_{d}^{iso}(\dot{x}_{i})=\eta_{iso}\dot{x}_{i}.$ (S.10)
where $x_{i}$ represents the position and node $i$ and $\dot{x}_{i}$ represents its velocity. Internal damping is selected to be proportional to the spring coefficients as [55]:
$F_{d}^{\text{int }}\left(x_{ij},\dot{x}_{i},\dot{x}_{j}\right)=\eta_{\text{int}}\left(\left(1-\frac{s_{ij}}{\left\|x_{ij}\right\|}\right)\left(\dot{x}_{i}-\dot{x}_{j}\right)+\frac{s_{ij}}{\left\|x_{ij}\right\|^{3}}\left[x_{ij}\cdot\left(\dot{x}_{i}-\dot{x}_{j}\right)\right]x_{ij}\right)$ (S.11)
where $s_{ij}$ is the rest length of the spring, and $x_{ij}=x_{i}-x_{j}$. $\eta_{iso}$ is selected to guarantee numerical stability, and $\eta_{\text{int }}$ coefficient is obtained by fitting the dynamic behavior to the experiments for different loading conditions. We obtained $\eta=0.05$ Ton s-1 by utilizing the Prony series characterized in [56].
A.3.3 Parameter Tuning
To determine each of our lumped model’s parameters, the DPF’s fundamental unit cell is isolated into two different parts (see Figure S5a). The dome unit is fitted as a nonlinear spring following Equation 1 and the strain limiting layer is represented by a liner spring to capture stretching and two rotational springs to capture the bending rigidity. Each of the constants for these springs needs to be tuned for different geometrical cases to produce an accurate model for inverse co-design. Given Equation 1, the parameters to be determined for the nonlinear spring are $k_{b}$, $\alpha$, and $d$. These parameters are functions of the dome unit’s mechanical behavior (stiffness and energy barrier) as well as the interaction between the dome and the wall of the air chamber of the subsequent unit cell. The procedure to find each constant as a function of the geometrical parameters of the dome unit cell is as follows. First, different finite element simulations of the unit cell are utilized to determine the best-fitted values for $k_{b}$, $\alpha$, and $d$ (Figure S5b) for different geometrical cases ($H$ and $t$, same as shown in Figure S1). Different parameters, such as elastic modulus, are examined to establish their influence on the strain energy and tip displacement (Figure S5b) and establish a better understanding of the governing parameters to be considered to represent the nonlinear spring accurately. Equation 1 is fitted to the FE data using the nonlinear least square method to obtain the overall behavior of these parameters in terms of $H$ and $t$. Results for these fitted parameters can be observed in Figure S5c, and the energy-displacement curve with the fitted parameters can be observed in Figure S6a.
Figure S5: a) DPF fundamental unit cell isolated into different lumped parameter spring elements. (nonlinear + linear + rotational springs) b) nonlinear spring parameter behavior over different dome heights $H$ and dome thickness $t$.
Once the values are obtained, we use Recursive Feature Elimination (RFE) [50] combined ridge regression to determine an expression for $k_{b}$, $\alpha$, and $d$ as functions of the dome height ($H$), dome thickness ($t$), and dome curvature ($R$). By using RFE, we can automatically determine the relevance of each feature in our model. Given the main task of inverse design of multistable actuators, a general and scalable model is needed to cover different design cases. We establish the model by considering three different non-dimensional relations:
$\pi_{1}=\frac{t}{H}\hskip 14.22636pt\pi_{2}=\frac{t}{R}\hskip 14.22636pt\pi_{3}=\frac{H}{R}$ (S.12)
where $\pi_{1}$ represents the load-carrying capacity of the shell, $\pi_{2}$ is the curvature-to-thickness ratio, and $\pi_{3}$ quantifies the dome shallowness. Using these three non-dimensional quantities, the following model is fitted:
[…]
$\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ is a matrix of all the possible candidates and interactions of the non-dimensional relations. The $N\times M$ matrix can be written as:
$\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})=\begin{bmatrix}\pi_{i}&\pi_{i}^{2}&\ldots&\pi^{n}_{i}&\pi_{i}\pi_{j}&\ldots&\pi^{n}_{i}\pi^{m}_{j}\end{bmatrix}$ (S.14)
where superscripts $n$ and $m$ represent the maximum degree of the polynomial use for the regression and subscripts $i$ and $j$ represent each of the non-dimensional numbers. For the parameter $\alpha$ we can fully represent this regression as:
$\begin{split}&\begin{bmatrix}\alpha(H_{1},t_{1},R_{1})\\ \alpha_{2}(H_{2},t_{2},R_{2})\\ \vdots\\ \alpha_{k}(H_{k},t_{k},R_{k})\end{bmatrix}=\\ &\begin{bmatrix}\pi_{i}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}(H_{1},t_{1},R_{1})&\pi_{i}\pi_{j}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{1},t_{1},R_{1})\\ \pi_{i}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}(H_{2},t_{2},R_{2})&\pi_{i}\pi_{j}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{2},t_{2},R_{2})\\ \vdots\\ \pi_{i}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}(H_{k},t_{k},R_{k})&\pi_{i}\pi_{j}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{k},t_{k},R_{k})\end{bmatrix}\begin{bmatrix}\xi_{1}\\ \xi_{2}\\ \vdots\\ \xi_{k}\end{bmatrix}\end{split}$ (S.15)
Figure S6: Parameter-fitting results for $\alpha$, $k_{b}$, $d$ as a function of geometric parameters. a) RFE with ridge regression cross-validation vs the number of features (training data) for $k_{b}$ (i), $\alpha$ (ii), and $d$ (iii); b) Prediction vs observation plot on test data for the reduced model for $k_{b}$ (i), $\alpha$ (ii), and $d$ (iii); c) Model fit for energy as a function of geometric parameters for monostable (i), metastable (ii) and bistable (iii) cases.
Note that we assume that the library of terms used to obtain $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ has a sufficiently rich column space that the behavior will be represented by Equation S.13, and it can be written as a linear combination of the weights $\xi$. The data is split into 70% train and 30% test data sets to calculate the accuracy of our model. Observation vs prediction plot and estimation error for the three different variables can be observed in Figure S6b (ii), where a good fit between the proposed model and the test data is observed. Moreover, we use RFE with cross-validation (cv) on the test data to eliminate features and retest the model to determine its accuracy. $R^{2}$ for the cross-validation data as a function of the number of features can be observed in Figure S6b(i). This gives a reduced model with similar prediction capabilities and the most relevant geometric relations of the problem, and yields an accurate representation of the behavior (see Figure S6b (iii)).
The final expressions for each of the nonlinear spring parameters can be written as:
$\begin{split}k_{b}=f\left(E,\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{E}{R^{2}}\Bigl{(}-\frac{1.97Rt^{5}}{H^{3}}+\frac{7.4H^{2}t^{3}}{R^{2}}+\frac{3.5Rt^{4}}{H^{2}}+\frac{0.37H^{2}t^{2}}{R}+\frac{42.2t^{5}}{H^{2}}\\ &-\frac{35.8Ht^{4}}{R^{2}}+\frac{71.8t^{5}}{HR}-\frac{3.4Rt^{3}}{H}-\frac{67.7t^{4}}{H}+4.2Rt^{2}+11.1t^{3}\Bigr{)}\end{split}$ (S.16)
$\begin{split}\alpha=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{0.29H^{3}}{R^{3}}-\frac{0.44t^{3}}{H^{3}}-\frac{2.2H^{2}t}{R^{3}}-\frac{0.73H^{2}}{R^{2}}-\frac{6.9t^{3}}{H^{2}R}+\frac{1.17t^{2}}{H^{2}}-\frac{3.0t^{3}}{HR^{2}}+\\ &\frac{7.05Ht}{R^{2}}+\frac{13.38t^{2}}{HR}+\frac{0.44H}{R}+\frac{13.67t^{3}}{R^{3}}-\frac{15.0t^{2}}{R^{2}}-\frac{4.7t}{R}\end{split}$ (S.17)
$d=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=2.14H+0.25t-H-\text{U}_{\text{sep}}-t_{ch}/2-t/2$ (S.18)
where $H$, $t$ and $R$ are dome unit parameters, $E$ is the elastic modulus of the material,$\text{U}_{\text{sep}}$ and $t_{ch}$ is the separation between units (see Figure S2).
A.4 Static and Dynamic Response
Using the parameters from the fitting process, we can predict stable state configurations for the DPF with the resulting model. Given that the nonlinear elements of the gripper are independently stable, the expected number of states for a given dome phalanx finger topology follows $2^{n}$, where $n$ is the number of bistable segments on the finger. The model can determine the behavior of a given design in orders of magnitude less time than the FE numerical simulations, which makes it feasible for iterating through potential configurations and optimizing for the best design according to a given task (e.g., position and grasping force).
Figure S7: Static and dynamic response of the DPF. a) Stable states are found by minimizing the strain energy. A comparison with FE simulations for five domes of the same height. b) Dynamic response of lattice model with viscoelastic dampers c) Schematic representation of applied load and final DPF position. d) Five-segment DPF dynamic response for different dome H.
Static response: The static response of the finger is predicted by minimizing the total energy of the geometry. We performed this process using Matlab’s fmincon sqp algorithm [57]. Our algorithm can predict different final shapes depending on the initial guess geometry. Results for the model are compared with FE simulations, where various points of the limiting layer are compared (see Figure S7a). Results show that our model can predict the final shape of the DPF for different geometric parameters with an error below 3% compared with FE simulations (see Table S1 for an extended comparison).
Dynamic response: The model can be further expanded by including point masses on each of the segments to represent the total mass of the system and dampers to capture the viscoelastic response of the material. By incorporating these elements and using symbolic differentiation, a dynamic model can be constructed to capture the time response of the DPF. The equations of motion for the system can be written as
$[M]\ddot{x}_{i}+F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})+F_{d}^{iso}(\dot{x}_{i},\dot{x}_{j})+\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)=F_{ext}(x_{i},t)$ (S.19)
where $[M]$ is the mass matrix of the system, $F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})$ is the internal damping, which can be modeled as a dashpot in parallel to each spring (see Figure 3a) and $F_{d}^{iso}(\dot{x}_{i})$ is the isotropic drag, which acts as an external force on each node proportional to velocity ($\dot{x}_{i}$[55]. $F_{ext}(x_{i,j},t)$ is the external actuation pressure that can be calculated as $F_{ext}(x_{ij},t)=A(x_{ij},)P(t)$. Where $A(x_{ij},)$ is the area of the face where the pressure is applied. By substituting $\dot{x}_{1}=x_{2}$ and $\dot{x}_{2}=\ddot{x}_{i}$, we can rewrite the system as
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -[M]^{-1}\left(F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})+F_{d}^{iso}(\dot{x}_{i},\dot{x}_{j})+\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)-F_{ext}(x_{ij},t)\right)\end{bmatrix}$ (S.20)
The system is solved using the Runge-Kutta time integration method for loading and unloading external force conditions. We explore the dynamic behavior of a five-segment DPF by applying the internal pressure as shown in Figure S7c, where we obtain the final stable state as shown and the dynamic response shown in Figure S7d. As expected, the model can capture the dynamic response of a bistable, metastable, and monostable DPF, where snap-through and snap-back phenomena are observed.
Furthermore, we analyze the temporal response of the metastable units to assess how geometrical parameters and material damping influence the resetting time. Our results show that adjusting the dome unit’s thickness modifies the resetting time (see Figure S9b). However, an even more pronounced effect is observed when varying the loading duration (see Figure S8c). This suggests that after manufacturing the DPF, additional tunability can be achieved by controlling the applied pressure (see Movie 7).
Figure S8: Lattice model dynamic response for dome unit cell. a) Internal pressure over time $P(t)$. Loading time $\tau_{1}$ and applied load time $\tau_{2}$. b) Effect of internal damping on dynamic response. c) Effect of applied load time ($\tau_{2}$) on metastable unit dome reseating time.
The resetting time response is validated against experimental data to assess its predictive accuracy. As shown in Figure S9c, the model effectively captures the time scales of a fully metastable DPF. However, due to the inherent randomness in the resetting process when all units are metastable, the tip displacement is not precisely predicted. Nevertheless, this tool enables the design of both the static and dynamic responses by programming the resetting time scales of the metastable domes (see Movie 2 for experimental comparison).
Figure S9: Metastable dome unit analysis. a) Effect of dome height in resetting time of the units (t = 0.8 mm). b) Effect of dome thickness in resetting time of the units (H = 3.0 mm). c) Comparison between experimentally recorded data shown in Figure 2c and our proposed model.
A.5 Model Validation and Simulation time
The lattice model is evaluated under different geometric configurations as shown in Table S1.
N Segments $R_{b}$ $H/R_{b}$ t E $t_{\text{lim}}$ Error %
1 5 5 0.7 0.7 5 1.4 1.37
2 5 5 0.8 0.8 5 1.6 0.62
3 5 5 1 1 26 2 0.51
4 5 8 0.7 0.75 5 1.5 0.51
5 5 8 0.8 1 26 1.6 2.73
6 6 5 0.7 0.7 5 1.4 1.42
7 6 5 0.8 0.8 5 1.6 0.6
8 6 5 1 1 26 2 0.4
9 6 8 0.7 0.75 5 1.5 0.44
10 6 8 0.8 1 26 1.6 2.47
Table S1: List of validation cases for spring model. Error % reference to the difference between model and FE analysis.
Different numbers of segments, elastic modulus, and limiting layer thickness are included to demonstrate the versatility of our approach. The error between the model and the simulation is calculated by comparing the difference between the final position of the limiting layer after all units are activated (see Fig S10).
Figure S10: Model comparison with FE method. (1)-(10) Case numbers parameters and error can be observed in Table S1.
Moreover, our model computational cost is evaluated by comparing the total simulation time simulation time with FE simulations. Results show that the model can yield the final shape of the DPF in less than 1 s. More importantly, the model does not exhibit a coupling between the number of segments and computational time. (see Table S2).
N segments FE Time [s] FE Time [h] Model Time [s]
2 3094 0.86 0.66
3 5960 1.66 0.58
4 12072 3.35 0.64
5 15859 4.41 0.76
6 21455 5.96 0.67
Table S2: Simulation time comparison between FE analysis and spring model.
A.6 Geometry Optimization
A.6.1 Inverse problem $\rightarrow$ Position Optimization
The final results for the position optimization can be observed in Figure S11, where the five target positions and the final shape predicted by the model are shown. The final geometry and objective function of the optimization algorithm are summarized in Table S3 and S4, where we can observe that for targets 1,2 and 5, the desired position is attained with almost perfect precision.
Figure S11: Final results for inverse design problem for tip position as an objective. a) Target 1, b) Target 2, c) Target 3, d) Target 4, e) Target 5, f) Objective function vs Number of segments for the first three target positions.
Table S3: Inverse design problem results for five different targeted positions (NinjaFlex 85A).
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.5 4.1 4.8 4.5 4.35
H2 4.7 4.3 4.8 4.7 4.35
H3 4.8 4.9 4.9 5.0 4.72
H4 5.0 4.9 5.0 4.87
H5 5.0 4.95
H6
Unit S 2.0 1.0 1.0 1.0 1.01
Unit L 9.9 7.0 7.3 7.0 8.99
t 0.6 0.62 0.68 0.62 0.75
$t_{\text{lim}}$ 1.5 1.3 1.2 1.4 1.25
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{dis}}$ 0.009 0.025 0.01 0.018 0.007
Obj 0.2 0.2 2.9 0.8 0.4
Table S4: Inverse design problem results for five different targeted positions (Cheetah 95A).
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.87 4.52 4.89 4.96 4.9
H2 4.8 3.93 4.72 4.52 4.87
H3 4.67 3.63 4.52 4.08 4.72
H4 3.98 4.33 3.99 4.35
H5 4.02 4.3
H6
Unit S 2.31 1.10 1.17 1.08 1.52
Unit L 9.83 7.28 6.95 6.93 8.8
t 0.79 0.83 0.65 0.76 0.63
$t_{\text{lim}}$ 1.63 1.59 1.19 1.07 1.72
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{dis}}$ 0.093 0.341 0.079 0.112 0.063
Obj 0.775 0.056 0.55 0.22 0.32
A.6.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
We further expand our model to include stiffness optimization in our objective function. By incorporating finger stiffness, grasping force is increased by changing the geometric parameters. Results for the optimization algorithm, where the same number of segments as in Table S3 and S4 is maintained, are shown in Table S5 and S6.
Table S5: Inverse design problem results for five different targeted positions with maximum stiffness (NinjaFlex 85A).
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.63 4.10 4.74 4.81 4.41
H2 4.67 4.27 4.83 4.82 4.58
H3 4.90 4.49 4.85 4.87 4.71
H4 4.98 4.86 4.92 4.81
H5 4.97 4.92
H6
Unit S 2.04 1.00 1.08 1.02 1.07
Unit L 9.94 7.00 7.07 7.03 8.85
t 0.83 0.86 0.81 0.85 0.83
$t_{\text{lim}}$ 1.5 1.36 1.45 1.35 1.4
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{st}}$ 0.011 0.0283 0.0146 0.0204 0.009
Improvement 1.22 1.13 1.46 1.13 1.3
Table S6: Inverse design problem results for five different targeted positions with maximum stiffness (Cheetah 95A).
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.99 4.66 4.96 5.0 4.9
H2 4.88 4.41 4.9 4.78 4.687
H3 4.86 4.31 4.72 4.69 4.57
H4 4.15 4.4 4.66 4.225
H5 3.92 4.2
H6
Unit S 2.26 1.4 1.11 1.01 1.03
Unit L 9.98 6.85 7.11 6.89 8.92
t 0.9 0.91 0.88 0.99 0.99
$t_{\text{lim}}$ 1.89 1.95 1.24 1.66 1.94
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{st}}$ 0.093 0.341 0.079 0.112 0.063
Improvement 1.51 1.61 1.48 2.7 2.37
A.6.3 Inverse problem $\rightarrow$ Position + Stiffness Optimization for multiple states
We design the multistable actuator by targeting three different tip positions. These targets are intended to be reached by activating 2,4 and 6 units. The following objective function is minimized:
$\begin{split}\min_{H_{i},t,\text{U}^{i}_{L},\text{U}^{i}_{\text{sep}}}&\,w_{1}\left(\text{T1}_{xy}-\text{Tip}_{\text{dis 1}}\right)^{2}+w_{2}\left(\text{T2}_{xy}-\text{Tip}_{\text{dis 2}}\right)^{2}\\ &+w_{3}\left(\text{T3}_{xy}-\text{Tip}_{\text{dis 3}}\right)^{2}+w_{4}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}\end{split}$ (S.21)
s.t. $\quad H_{i+1}\geq H_{i},\quad i=1,\dots,N-1,$
$\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N,$
$\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N.$
Table S7: Experimental measurements for the multistable actuator.
Target Coordinates Model Experiments
State 1 ($T_{1}$) [54 , 40] [52.95 , 32.82] [ 50.2, 30.6]
State 2 ($T_{2}$) [40 , 45] [40.84 , 45.9] [ 38.2, 44.2]
State 3 ($T_{3}$) [30 , 45] [33.81 , 47.35] [ 35.2, 46.1]
Where $w_{1}=w_{2}=w_{3}=w_{4}=1/4$ and $T_{1}$ to $T_{3}$ represent the target positions for the gripper as shown in Table S7. Results for each of the targets can be seen in Figure S12.
Figure S12: Inverse optimization results for multistable dome phalanx actuator. a) Lattice model results for $T_{1}$(i), $T_{2}$(ii) and $T_{3}$(iii). b) Experimental results for $T_{1}$(i), $T_{2}$(ii) and $T_{3}$(iii).
A.6.4 Inverse problem $\rightarrow$ Bistable + Mestastable units
Our task planning application uses our dynamic model to calculate the resetting time of the metastable units. We fix our actuator geometry by selecting six segments with four bistable and two metastable units. The selected geometry parameters are $H_{1}=5$ mm, $H_{2}=5$ mm and $H_{3}=5$ mm, $\text{U}_{\text{L}}=6.5$ mm and $\text{U}_{\text{sep}}=1$ mm, $t=0.8$ mm, and the last two-unit heights ($H_{5}$ and $H_{6}$) is designed to maximize the resetting time under the two different pressure profiles show in Figure S13a. Both dome heights are constrained to be the same ($H_{5}=H_{6}$), and the initial bistable set point can be observed in Figure S13b. The full dynamic response predicted by the model is shown in Figure S13c, where we can observe the three different snap-through instability and two snap-backs in the two simulated cases. A restriction is imposed over these units to guarantee that they were metastable, by using the state map reported Figure S1b. As expected, to maximize the resetting time at two different conditions, the algorithm would converge to the highest possible metastable unit (see Figure S1b).
Figure S13: Pick and place task planning set points and model inputs. a) Pressure profile for task planning. b) Setpoint (stable state) for the first active three bistable units. c) Dynamic response for the bistable + metastable actuator.
A.7 Experimental Test
A.7.1 Pressure and displacements Measurements
As shown in Figure S14, an air compressor (Pump DOA-P704-AA) is used to supply air to the DPF. The pressure is measured using a HONEYWELL ABPDANN010BG2A3 pressure sensor (0-10 Bar range), which is connected to an Arduino Uno for data acquisition. The Arduino records pressure readings at 500ms intervals to ensure accurate monitoring. The pressure is gradually increased to a specified value before being released to invert the dome units. The dynamic behavior of the DPF is captured in detail using a Photron Fastcam Mini UX100 high-speed camera, as illustrated in Movie 1. Tip displacement over time is then analyzed using the Tracker Video Analysis and Modeling Tool, providing precise measurements of the system’s response.
Figure S14: Experimental setup for pressure and displacement measurement. Schematic of the test setup used to record the displacement and pressure over time of the DPF and DFR. (1) DPR or DPF (2) Air Pump. (3) Pressure sensor. (4) Arduino Uno. (5) Laptop. (6) High Photron Speed camera.
A.7.2 Inverse problem experimental validation
Five geometries are tested by activating all bistable unit cells and measuring the final tip position. The final position of the finger is measured by using image processing. Results can be observed in Figure S15a (i-v) for NinjaFlex 85A and Figure S15b (i-iii) for Cheetah 95A, and the target coordinates, model prediction, and experimental measurements are reported in Table S9
Figure S15: Experimental results for different DPF geometries. a) Printed samples for inverse design target positions (i) - (v) for Ninjaflex 85A b) Printed samples for inverse design target positions (i) - (iii) for Cheetah 95A.
Target Coordinates Model Experiments Error
[46 , 32] [46.85 , 31.22] [42.9 , 37.7] 13%
[25 , 20] [25.34 , 20.34] [23.53 , 22.93] 10%
[25 , 40] [30.10 , 40.70] [29.4 , 43.5] 6%
[25 , 30] [28.76 , 29.87] [23.0 , 34.26] 17%
[35 , 45] [37.98 , 44.36] [36.8 , 50.77] 11%
Table S8: Experimental measurements for five different DPFs optimized for a specific target position (NinjaFlex 85A).
Target Coordinates Model Experiments Error
[46 , 32] [ 45.6, 31.8 ] [43.31 , 28.38] 6.9%
[25 , 20] [ 24.7, 19.5] [ 25,17, 16.55] 4.3%
[25 , 40] [ 25.15, 39.3] [ 23.03, 39.29] 2.5%
[25 , 30] [25.2, 29.22] [ 22.9, 31.6] 1.1%
[35 , 45] [ 33.97, 45.3] [ 33.35, 44.9] 1.2%
Table S9: Experimental measurements for five different DPFs optimized for a specific target position (Cheetah 95A).
A.7.3 Experimental validation stiffness optimization
The stiffness of each DPF is measured using an Instron universal testing machine and applying a perpendicular force to the tip of the actuator (see Figure S15b and Movie 5). Linear behavior is observed in the neighborhood of each stable state (0, 2, 4, and 5 active units in Figure S15b), with a different stiffness for every state. Finally, the inverse design problem results are tested by comparing the stiffness of two geometries with the same target position but two distinct objective functions (Equation 6 and Equation 8).
Figure S16: Force vs displacement plot for different DPF. a) Experimental setup. b) Effect of number of active domes on DPF stiffness. c) Comparison between DPF geometries for position inverse design and position+stiffness inverse design for target 5 (Cheetah 95A and NinjaFlex 85A).
A.7.4 Durability Test
The robustness of the DPF design is tested by performing 100 cycles over three weeks to show the geometry’s repeatability and effect. Results show that after three weeks and 100 cycles per week, the tip position stays within 5% of the target tip displacement.
Figure S17: Durability test and time decay response for a) NinjaTek NinjaFlex b) NinjaTek Cheetah (95A).
Notice that the characteristics of multistability reduce the potential effects of material variability by making geometry (i.e., morphology) the dominant driver of our soft robot’s response.
A.7.5 Carrying Capacity
Payload capacity and object grasping tests were conducted to evaluate the performance of our DPR, demonstrate its grasping capabilities, and determine the maximum carrying weight of the system. The designed global states enable adjustments in output force and position based on the number of inverted domes (see Movie 6). This adaptability allows the DPR to grasp objects of various sizes and topologies by removing the actuation pressure. Figure S18 illustrates the range of objects successfully grasped by the DPR using only the energy stored in the structure due to dome inversion. As depicted, diverse object types are effectively grasped, highlighting the system’s versatility. Additionally, to characterize the payload capacity, weights were incrementally added to the object shown in Figure S18 (i) until the structure could no longer grasp it. A maximum load of 738 g was recorded, resulting in a maximum load-to-weight ratio of 10.25 (See Movie 6).
Figure S18: DPF grasping capabilities are shown with six different objects. a) Blackberry Replica, b) Tennis Ball, c) Toy Dinosaur, d) Baking Soda pack, e) Mini Fruit Can. f) Plush Toy g) Cylindrical Cream. h) Rubik’s Cube. i) Coffee Cup.
A.7.6 Scalability
As the mechanical response of the DPF is derived from the dome geometry and its stability, the DPF and robot can be scaled to increase to large dimensions just by considering the energy stored during the dome inversion. Following the analysis by Seffen and Vidoli [58], we can derive different additional relations for scaling, given the energy required for inverting a single dome. This energy is proportional to $\frac{Et^{3}}{1-\nu^{2}}\left(\frac{H}{R}\right)^{2}$ where E, $\nu$, t, and R are the elastic modulus, Poisson’s ratio, thickness, height, and radius of the domes respectively.
We can utilize the dome shallowness $\frac{H}{R}$ and the curvature-to-thickness ratio $\frac{t}{R}$ to either scale up or down the DPF geometry. To test this, we 3D printed two different geometries with the same energetics but a 2x difference.
Figure S19: DPF scalability test. a) Two different DPF specimens scale using adimensional relationships. b) Stable states with all dome units active and a comparison with our lattice model.